A Bibliography of Publications in Theoretical Computer Science: 1985–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

08 February 2014 Version 1.12

Title word cross-reference

2 [214, 81, 498]. α_{0} [468, 227, 256]. α_{2} [256]. $beta [546]. β [395]. D′_{1}∗ [255]. exp(Θ( ^{p} \sqrt{n})) [414]. F [401]. G[u]/ < Q(u)^{ℓ} >, ℓ > 1 [424]. γ [491]. Gw [577]. k [65]. l [265]. $lambda [559, 526]. Λ [91, 395]. N^{k} [131]. n^{th} [26]. n^{Θ((ln n)^{(}1/p))} [414]. O(modT mod ^{3}) [7]. O(k.2^{n}k) [310]. O(n^{lgk}.2^{n}2) [310]. ω [477]. Ω(n^{4}/3) [26]. X [452].

-acyclic [491]. -ary [65]. -calculi [559]. -calculus [395]. -colourability [81]. -context-free [498]. -expansions [546]. -groups [265]. -languages [477, 577]. -product [468, 227]. -products [256]. -programs [91]. -regular [469]. -terms [526]. -trees [452].

14th [640].

2-asynchronous [474].

’86 [635]. ’87 [637].

Abelian [11, 189, 133, 394, 265]. Abstract [207, 74, 448, 630, 57]. Abstraction [563, 306, 51, 543]. Accepting [477]. acceptors [579, 164]. access [423]. Ackermann [422]. actions [345]. active [270]. acyclic [10, 491, 479, 439]. adaptive [569]. Adding [59]. addition [214, 540]. adherences [213]. Adic [556]. adjacent [141]. Adjunction [134]. advantages [408]. Alexander [628]. Algebra [339, 51, 424, 311, 36, 198]. algebra-framework [198]. Algebraic [296, 127, 629, 206, 305, 166, 112, 222, 167, 285, 411, 364, 151, 575, 20, 57, 415, 154]. algebras [515, 91, 405, 136]. algorithm [371, 190, 131, 470, 7, 609, 291, 302, 180, 439, 410, 522, 442, 310, 565]. algorithmic [272, 222, 117]. Algorithms [479, 139, 423, 424, 36, 287, 43, 529, 331, 623, 531, 348, 109, 399, 365]. ALNA [118]. alphabet [93, 136]. Alternating [152, 478, 42, 373, 149, 511, 592, 579, 378, 590]. Alternation [164]. ambiguity [292]. among [95]. Amortized [273]. analysing [222]. Analysis [302, 481, 600, 40, 607, 291, 556, 161, 497]. Analytic [292]. anonymous [346]. answer [403]. aperiodic [536]. application [383, 17]. Applications [55, 284, 418, 153, 527, 406, 289, 266]. approach [206, 423, 76, 207, 388, 621, 408, 628, 587, 575, 92, 563, 109]. Approximating [267]. Approximation [365, 272, 561, 410]. approximative [226, 561]. arbiters [175]. arbitrary [290]. arbre [463]. arithmetic [45, 400, 556]. arithmetical [528]. Arithmetics [544, 639]. arrays [452, 416, 406]. ary [65]. aspects [295, 525]. assertion [447]. assignment [218]. assignments [202, 118, 27]. associated [367]. Asynchronous [374, 474, 269, 285, 77]. asynchrony [58]. ATMs [604]. atomicity [444]. attribute [341]. Automata [282, 634, 640, 407, 115, 419, 264, 338, 464, 185, 240, 268, 474, 269, 270, 227, 350, 392, 355, 129, 156, 608, 145, 579, 478, 636, 69, 42, 477, 590, 50, 373, 8, 183, 512, 177, 157, 289, 149, 98, 234]. automated [221]. Automatic [91, 359, 467]. automation [589, 108]. automaton [508]. aux [549]. Average [219, 497]. average-case [497]. Avoidable [502, 633]. axiomatic [383, 369]. Axiomatisation [6]. axiomatization [385]. backtracking [607]. Balance [248, 603]. Bandwidth [140, 612]. based [465, 207, 74, 564, 274, 525, 624, 317]. baseline [533]. baseline-equivalent [533]. bases [159, 139]. basis [348]. BCK [526]. BCK-combinators [526]. BCNF [491]. be [144]. behavior [632]. behavioral [512]. behaviour [382, 374, 521]. behavioural [503]. belief [431]. benchmarks [532]. Bendix [571]. Bergstra [591]. beta [582]. between [207, 425, 288, 523, 391, 459, 239]. biconnected [516]. Bilinear [486, 424]. binary [44, 574]. biology [527]. bipartite [351]. biprefix [428]. bisimulation [357]. bitonic [221]. blocking [534]. Boolean [352, 122, 159, 69, 570, 32]. bornée [463]. bound [26, 125]. boundary [270]. Bounded [39, 313, 115, 119, 338, 604, 456, 318, 378, 501, 46, 519]. Bounded-depth [39]. bounded-width [501]. Boundedness [183]. bounds [38, 572, 226, 253, 214, 152, 377, 409, 529, 333, 64, 520]. bracket [354]. branching [539, 217, 501, 522]. broadcast [331]. broadcasting [287]. Brooks [609]. Brzozowski [277]. bubbling [376]. Buchi [234, 289]. Building [344]. Burnside [107].

c [582, 392]. C-tree [392]. Calculation [5, 243]. calculi [127, 559, 399]. calculus [105, 134, 596, 395, 624, 321]. cancellation [83]. Canonical [375, 300, 21]. case [607, 497, 337]. cases [513]. Categorical [295, 559, 134, 340, 558]. Categories [617]. CCC [523]. CCL [223, 559]. CCS [615]. cellular [145, 416]. central [194]. Certain [619, 184, 597, 367, 310]. CFL [284]. chain [618, 35, 175]. chains [5]. change [386]. channel [331, 183]. character [73]. characterisation [305]. characteristic [43]. Characteristics [304]. Characterization [156, 212, 533, 3, 202, 357, 30, 16, 492, 67, 590, 191, 308, 158, 46, 85, 369]. Characterizations [2, 334, 406, 213, 581, 329, 316].

Characterizing [446]. characters [73]. checking [522]. choice [389]. Chomsky [30]. chordal [351]. Church [105, 7, 90, 623, 280, 278, 565, 558]. circuits [39, 409, 77]. Circular [419, 517]. class [334]. classes [504, 468, 463, 227, 429, 19, 614, 597, 71, 2]. Classical [341]. Classification [424, 547, 37, 238]. clausal [541]. clause [616]. clauses [186, 454]. closed [534]. closure [121, 416, 484, 173, 439]. closures [303]. Clustering [78]. CNF [612]. co [284]. co-CFL [284]. coarsest [110]. Code [18, 68, 618, 390, 35]. Code-compatible [68]. Coded [181]. codes [132, 419, 619, 29, 143, 106, 367, 509]. Coding [544, 639]. coefficients [424, 180]. Cohesion [505]. Colloquium [339, 282, 640, 636]. colourability [81]. colouring [178]. combinator [388, 148, 568]. combinatorial [423, 132, 168, 398, 518]. combinators [105, 526, 558]. combinatory [482, 582, 563]. Comment [626]. Comments [338]. common [410]. communicating [86, 51, 285, 437, 183, 293]. Communication [263, 525, 199, 253, 408]. communications [286]. commutations [524]. Commutative [225, 404, 216, 417, 276, 434, 136]. commutativity [586]. compactness [418]. Comparison [423, 472, 569, 372]. comparisons [179]. compatibility [308]. compatible [68]. complementation [318, 593, 234, 289]. Complete [189, 227, 169, 599, 15, 504, 202, 225, 513, 429, 48, 529, 435, 327, 229, 560, 369, 219, 140, 310]. completeness [167, 218, 94, 22, 543, 372, 351]. completing [101]. Completion [598, 571, 602]. Complex [32, 180]. Complexity [600, 594, 429, 148, 14, 362, 35, 398, 407, 504, 346, 319, 26, 475, 259, 219, 194, 178, 47, 226, 214, 263, 408, 394, 41, 409, 614, 554, 238, 146, 438, 252, 71, 328, 150, 520, 519, 459, 567, 63, 170, 265]. complicated [303]. composition [180]. compositions [135]. compositive [563]. Comput [634]. Computability [218]. computable [307, 296, 251, 379]. computation [465, 352, 604, 237, 122, 422, 151, 344]. computational [319, 394, 182]. computations [23, 620, 362, 89, 455]. compute [235]. computed [570]. Computer [82, 635]. computers [159]. Computing [133, 424, 345, 376]. Concatenation [508]. conceptual [391]. Concerning [141, 254, 214, 538, 564]. Concurrency [444, 199, 283]. Concurrent [171, 537, 380, 535, 423, 382, 74, 571, 25, 629, 433, 89, 438, 440]. condition [573, 584, 144]. conditional [128]. conditions [209, 420, 477, 340]. cones [411]. Conference [82, 638, 637, 544, 639]. configurations [396]. conflict [538, 492]. conflict-free [538, 492]. Confluence [559, 586]. congruence [104, 598, 420]. congruences [36]. conjecture [138, 591, 93, 224, 198]. conjugacy [14, 229]. conjugaison [463]. conjugation [463]. connection [239]. connections [533]. consequence [254]. consistency [306]. constrained [140]. constraints [300]. construct [139]. constructed [547]. constructibility [13]. constructing [36, 470, 410]. Construction [29, 367, 509, 476, 159, 615, 124, 192]. Constructive [311]. constructor [473]. constructors [295]. contact [409]. containment [490, 163, 358]. contains [115, 338]. Context [83, 115, 402, 338, 130, 383, 141, 59, 292, 30, 28, 288, 135, 191, 128, 414, 252, 498, 236, 316]. Context-free [83, 402, 130, 383, 141, 59, 292, 30, 28, 135, 414, 252, 236, 316]. context-sensitive [115, 338, 191]. Continuation [197]. continued [251, 379, 552]. Continuous [176, 160].

Continuously [622]. contribution [53]. control [74, 332]. controlling [423]. controls [74]. convolution [26]. coordinated [49]. core [238]. cores [475]. correctness [299]. correspond [196]. countable [261]. counterexample [105]. Counting [427, 47, 494, 239]. covering [371]. coverings [80]. covers [492]. creation [460]. creative [242]. CREW [625]. Critical [468, 302]. cryptosystem [564, 274]. CSP [286, 374]. CTS [114]. cubic [516]. cut [407, 435]. cycle [38]. cycles [476, 190]. Cyclic [355, 229].

D0L [18, 211, 262]. D0L-systems [262]. daisy [175]. DAL [40]. DAL-a [40]. dans [466]. data [296, 207, 517, 40, 295, 273, 20, 300]. database [489, 423, 10, 638, 491]. databases [364, 393, 336]. Deadlock [451, 328]. decentralized [175]. Decidability [596, 276, 35]. decidable [232, 21, 280, 144]. decide [131, 565]. Deciding [81, 586, 438]. decision [370, 64]. Declarative [632]. Decomposition [157, 208, 120, 250, 89]. Decompositions [432, 561]. definability [485]. definable [321]. defined [360, 186]. defining [408]. definite [616]. definition [383]. degree [26]. degrees [314, 579]. degrees-rotation [579]. Dehn [302]. Delaying [399]. demonstration [162]. Dense [184, 476, 178]. dependencies [66, 101, 147, 364, 492]. dependency [97, 495]. dependent [288]. depth [39, 441, 576]. ’Deque [338, 115]. Derivation [345, 113, 201, 196, 279]. derivations [415]. deriving [177]. description [618]. descriptive [597]. Design [532, 54]. Designing [460]. desirability [491]. detecting [233]. detection [38, 183, 328]. Determinacy [24]. determinant [555]. determination [387]. Determinism [482]. Deterministic [579, 166, 264, 335, 28, 151, 522]. development [637]. devices [56]. diagonal [552]. diagonalization [32]. Diagonalizations [307]. dictionary [425]. different [109]. digital [430]. digraphs [479, 439]. dimension [290]. dimensional [281, 214, 406, 592, 579]. direct [522]. disambiguating [506]. discipline [449]. Discrete [370, 319]. Disjunctive [208]. distance [78]. distinction [207]. distinctness [247]. Distributed [287, 331, 286, 529, 291, 616]. distribution [168]. divergent [602]. divisibility [15]. DL [313]. Domain [336]. Domain-independent [336]. domains [605, 360]. dominance [358]. dominating [487, 351]. domination [368]. dot [441, 576]. dot-depth [441, 576]. down [270]. DPDAs [144]. d’un [466]. d’une [463]. Dyck [255]. dynamic [527, 220, 452, 455, 27]. Dynamics [129, 546].

E-CCC [523]. E-unification [599]. ear [250]. easiest [45]. easy [533, 235, 233]. edge [304, 435]. EDS [250]. effective [206]. efficiency [273]. Efficient [604, 172, 88, 345, 455, 442, 565]. Ehrenfeucht [138, 224, 198, 93]. election [361]. element [247, 143, 106]. elements [312, 325]. Eleventh [339]. eliminants [121]. elimination [531]. embeddable [261]. embedding [452, 246, 245]. Embeddings [44, 617]. empty [183]. ends [50]. engendered [619]. entailment [391]. Entire [93]. Entropie [595, 483]. Entropy [409, 419, 595, 483]. enumerable [496, 16]. equality [384, 229, 186, 31]. equation [615]. Equational [517, 405, 385, 169, 480, 596, 543]. equations [206, 197, 216, 514, 87, 103, 187, 93, 568]. equivalence [349, 3, 232, 24, 230, 330, 288, 433, 135, 248, 267, 193, 512, 85, 144, 522, 72]. equivalence=trace [24]. Equivalences [153, 503]. equivalent [460, 533, 579, 21].

EREW [625]. ergodic [552]. Errata [634]. error [510]. errors [207]. Etude [436]. Euclidian [608]. Euler [553]. evaluating [376, 150]. evaluation [137, 341, 167]. Event [605]. events [2]. Every [261]. evolution [171]. Exact [534, 344]. exception [207]. exceptions [207]. existence [10, 57]. expansions [546]. expressions [264, 217, 127, 75, 19, 182, 175]. expressiveness [94]. extended [607, 628, 175]. Extending [75]. extension [105, 20, 326]. Extensional [445]. external [603]. Extraction [481]. extremal [117].

F [202, 203]. F-complete [202]. facteurs [466]. factor [420]. factorial [589]. Factoring [325]. factorisation [29]. factors [573, 532, 258, 466]. failure [560]. fair [306, 58, 285, 438, 583]. fairness [451, 538, 366]. families [384, 97, 503]. family [367, 509]. Fast [588, 36, 145, 43, 89, 180]. Fibonacci [528]. FIFO [12]. fifteen [203]. figures [4]. file [497]. Finding [4, 41, 487, 528]. fini [466]. Finite [240, 634, 428, 165, 419, 446, 403, 163, 232, 36, 214, 97, 579, 506, 21, 478, 597, 42, 229, 623, 246, 586, 8, 367, 183, 438, 93, 466, 283, 509]. finite-state [438]. Finitely [557, 209]. firing [297]. First [272, 638, 254, 369, 421, 473]. First-order [272, 254, 369, 421, 473]. Fixed [515, 485, 622, 204, 195]. fixed-point [204]. fixed-points [195]. fixedpoints [84]. fixpoint [514, 317]. flowchart [337]. flowgraph [197]. flows [335]. Floyd [22]. fonction [549]. forbidden [117]. forbidding [59]. forests [201]. form [541, 582, 124, 192]. Formal [66, 481, 315, 55, 136]. forms [137, 375]. formulas [612, 611, 79, 570, 336, 212]. Forys [626]. Foundations [82, 635]. Fourier [319, 588]. fractal [549]. fractales [549]. fraction [251, 379, 552]. fragment [624]. framework [506, 198]. Fredholm [555]. free [515, 402, 130, 383, 141, 225, 216, 59, 292, 83, 276, 30, 538, 28, 188, 492, 67, 284, 135, 5, 434, 414, 252, 498, 57, 236, 316, 174, 495]. freedom [492]. French [549, 463, 436, 595, 466]. Full [543, 270]. function [45, 422, 584, 110, 376]. functional [514, 101, 97, 610, 364]. Functions [570, 545, 352, 427, 549, 194, 39, 375, 159, 580, 96, 249, 13, 519, 63, 170]. Further [430].

Galois [235]. gate [471]. Gaussian [343, 531]. general [319, 599, 133, 252]. generalisation [70, 68]. generalised [34]. Generalization [507, 467, 440]. Generalized [217, 187, 475, 384, 500, 588, 353, 494]. generated [622, 113, 304, 209, 602, 557]. Generating [412, 190, 618]. Generation [580, 168, 390, 463, 463]. generator [41]. generators [305, 347]. Gentzen [567]. Geometric [363, 358, 4, 525]. geometrical [587]. geometry [608, 553]. German [483]. Girard [401]. given [258]. global [59]. grammar [326]. Grammars [279, 153, 383, 59, 270, 456, 83, 196, 245, 128, 49, 381, 99, 236, 566]. grammatical [590, 80]. graph [383, 428, 456, 304, 323, 488, 196, 472]. graph-rewrite [472]. graphs [368, 476, 178, 470, 614, 516, 250, 351]. greatest [195]. greedy [410]. Greibach [124, 192]. Gries [440]. ground [280]. group [235, 394, 398, 565]. groups [189, 133, 367, 265]. Growth [633, 1].

H [81]. H-resolution [81]. Hamiltonian [476]. handling [207]. hard [504, 133]. having [492]. HDT0L [232]. helping [322]. Herbrand [23]. Herbrand-Kleene [23]. Hermite [139]. Heuristics [574]. Hierarchical [77, 489]. Hierarchies [268]. hierarchy [363, 611, 400, 263, 592, 116, 348, 277]. high [205]. high-level [205]. Higher [56, 405, 507, 218, 381]. Higher-level [56]. higher-order [405, 507]. Higman [70]. histories [275, 505]. Hoare [621, 94, 440]. homeomorphism [76, 614]. homeomorphs [47]. Homogeneous [120]. homomorphic [384, 30, 16, 191]. homomorphisms [496, 18]. Honest [314]. Horn [611, 186, 317]. hypergraph [81].

Ianov [342]. ideals [36, 606]. idempotent [404]. identifying [73]. identities [553]. II [69]. II. [335, 337]. implementation [286]. Implementing [473]. Implication [454]. Implications [117]. implicit [425]. impossibility [191]. Improved [38, 439]. incidence [525]. Incompleteness [421]. incremental [341]. Independent [533, 156, 470, 336]. index [255, 414]. indexed [245, 279]. individual [142]. induction [360, 173]. Inductive [447, 359, 161]. inefficient [513]. inevitability [537]. inference [490, 585]. Inferences [147]. infinis [436]. Infinite [554, 160, 283, 184, 374, 70, 215, 602, 257, 284, 329, 373, 436, 177, 99]. Infinite-word [160]. infinity [469]. inherently [422]. initial [207, 262]. inputs [508]. institutions [57]. integers [174]. Integrity [136]. intensional [357]. Interaction [495]. Interaction-free [495]. interchange [402]. intercluster [78]. interconnection [533, 260]. interleaving [521]. International [282, 635, 637, 640, 638, 636]. interpolation [150]. interpretation [401, 630]. Interpretations [450]. intersection [426, 394, 291, 492, 449]. interval [375]. intractability [100]. Introduction [493, 311, 12, 366]. invariant [453]. Inverse [576, 135]. inversion [453]. inverted [497]. invertible [580]. investigation [74]. irrational [551]. isolated [407]. isomorphism [614, 516, 95]. isomorphisms [249]. ITA [228]. iterated [60, 75, 564, 274]. iterations [266]. iterative [422, 406].

join [66]. joint [637].

kernel [154]. key [564, 274]. Kleene [23]. Klop [105]. Knowledge [431, 575]. known [510]. Knuth [571]. Koomen [306]. Kripke [446].

labelled [548]. Lambda [134, 105, 182, 596, 399, 321]. lambda-calculus [105, 321]. langage [466]. Language [346, 577, 589, 338, 384, 217, 255, 55, 584, 511, 145, 187, 466, 440, 398, 154]. Language-theoretical [577]. Languages [282, 634, 640, 104, 115, 315, 184, 402, 338, 74, 240, 268, 232, 141, 632, 292, 496, 208, 618, 304, 429, 209, 420, 211, 313, 30, 16, 28, 610, 37, 330, 636, 484, 69, 347, 135, 469, 329, 477, 191, 279, 92, 436, 414, 160, 107, 31, 46, 252, 354, 324, 498, 441, 35, 195, 277, 543, 212, 577, 316, 566]. largest [357]. later [203]. lattice [139, 348]. lattices [238]. laws [461]. lazy [610]. LC [566]. leader [361]. Leaf [592]. Leaf-size [592]. Lech [162]. Left [606]. lemma [28]. lemmas [402]. length [463, 41, 603, 5, 397]. letters [502]. level [111, 205, 390, 630, 381, 56]. Lie [231]. like [387]. limited [463]. limiting [543]. line [152, 579]. linéaire [466]. Linear [294, 413, 228, 462, 281, 526, 179, 87, 448, 554, 555, 110, 157, 64, 624, 466, 369]. linear-time [228]. lines [44]. list [343]. Literal [315]. local [27]. Löf [137, 301]. log [377]. log-space [377]. logarithmic [281]. logic [413, 613, 517, 447, 632, 40, 271, 514, 294, 220, 559, 79, 621, 538, 215, 628, 84, 393, 118, 582, 77, 50, 327, 563, 94, 494, 289, 624, 440, 254, 369, 421, 27, 567, 627, 446]. Logical [485, 3, 92]. logico [575]. logico-algebraic [575]. logics [380, 299, 293, 459]. longueur [463]. looping [482]. loops [217]. Lower [572, 226, 253, 64, 520, 38, 26, 377, 409, 125, 529, 333]. lower-bounds [333]. LRR [326]. Lyndon [463].

machine [148, 116, 448, 386, 312, 46]. machines [119, 155, 429, 606, 318, 511, 152, 592, 322, 378, 593, 49]. macro [155]. Mahler [162]. Manipulating [201]. manipulation [530]. manipulations [370]. many [248, 327]. many-valued [248, 327]. mappings [160]. Martin [137, 301]. Martin-Löf [137, 301]. MAT [497]. matchers [169]. Matching [462, 320, 172, 365]. matchings [476]. mathematical [34, 635]. matrices [530]. Matrix [411, 453, 362, 17]. maxima [303]. Maximal [60, 403, 470, 509]. maximally [171]. maximum [78]. means [341, 204]. measures [259, 550]. medium [98]. MEIJE [56]. MEIJE-SCCS [56]. membership [394, 330]. memory [270]. Merging [204]. method [222, 447, 628, 89, 494, 22]. MFCS [635]. Min [435]. mincing [486]. minima [303]. Minimal [589, 424, 36, 297]. minimize [78]. minimum [41, 574, 487, 442]. minimum-length [41]. Minkowski [139, 552]. mismatches [320, 172]. Mixed [269, 534]. mn [371]. Modal [541, 293, 393]. modality [33]. model [575, 437]. modeling [632]. Models [333, 460, 445, 604, 271, 292, 625, 89, 57]. modes [279]. modular [86]. modulo [424, 598, 586]. molecular [527]. monadic [623, 212]. monogeneous [12]. monoid [565]. monoide [5]. monoids [404, 309, 11, 225, 216, 276, 434, 14, 278, 536, 576]. monotone [352, 26, 562, 570]. morphic [316]. morphisms [451, 564, 67, 135, 274]. Morse [518]. most [387]. most-recent [387]. mots [463, 436]. multi [462, 331]. multi-channel [331]. multi-linear [462]. multicounter [318, 593]. Multidimensional [396]. multihead [268, 478, 42]. multiple [187]. Multiplication [45, 125]. multiplicity [31]. multipushdown [46]. multiset [587]. multistage [533]. multivalued [364, 492, 495]. MVD [492]. MVD-intersection [492].

n [111, 312]. n-level [111]. name [102]. name-stamp [102]. Narrowing [443]. natural [48, 393]. NC [609]. nearly [470]. necessary [584]. Neciporuk [539]. necklace [573]. negation [560]. nested [493]. Net [196, 382, 127, 92]. nets [114, 163, 538, 205, 524, 12, 142, 521]. nets-monogeneous [12]. network [26, 529, 260]. networks [534, 533, 54, 562, 122, 287, 129, 253, 331, 512, 32, 525]. NLC [383, 456]. no [570]. Non [97, 225]. non-commutative [225]. Non-finite [97]. Nonapplicable [210]. noncomplete [96]. nondeterminacy [450]. nondeterminism [199, 237, 501]. Nondeterministic [593, 166, 23, 621, 318, 389, 116, 432, 118, 337]. nonextensional [134]. nonmonotone [562]. nonpolynomial [96]. nonsequential [171]. nonstandard [621, 299]. Nonterminal [456]. nontrivial [45]. normal [582, 124, 192]. note [305]. Notes [209, 420]. notion [257]. notions [372]. NP [219, 513, 242, 314, 96, 140, 435, 351, 327, 233, 310, 303]. NP-complete [219, 140, 310, 513, 435, 327]. NP-completeness [351]. nulls [210]. number [528]. numbering [250]. numbers [251, 379, 528]. numeration [549, 549]. numerical [147, 180].

O [371]. object [275, 505]. objects [489]. Observation [349, 3, 24]. observational [85]. Observations [564, 239, 283]. obtained [32, 99]. occam [461]. occurrence [548, 258]. off [179]. omega [104, 167, 37, 164, 347, 218, 173, 436, 175, 195, 277, 212]. omega-completeness [167]. omega-extended [175]. omega-languages [104, 436, 277]. omega-power [347]. omega-regular [37, 212]. omega-sups [173]. omega-tree [195]. omega-type [164]. on-line [152, 579]. One [395, 268, 249, 116, 522, 98, 72]. one-one [72]. One-step [395]. one-tape [116]. one-way [268, 249]. only [152]. open [534]. operation [321]. operational [632]. operations [384, 528]. operator [389]. operators [48]. Optimal [452, 529, 84, 531, 354, 573, 470]. optimistic [471]. optimization [363]. Optimizers [499]. optimum [574]. oracle [322]. oracles [15, 109]. orbit [184]. order [405, 272, 507, 561, 581, 50, 254, 369, 421, 473, 212]. ordered [105, 412]. ordering [126]. orderings [587]. oriented [9]. other [368, 105]. outerplanar [516]. overlap [253]. Owicki [440]. Owicki-Gries [440].

P [218]. P= [314]. packet [353]. pair [105, 144]. pairs [302]. palutators [398]. papers [638]. Paraconsistent [613]. Parallel [476, 122, 320, 250, 604, 237, 159, 388, 470, 145, 530, 614, 532, 362, 312, 531, 252, 354, 455, 328, 266, 442]. parallelism [199]. Parameter [381]. Parameter-reduction [381]. parameterized [308, 300]. parametrization [507]. parsing [252, 354, 326]. part [515, 424]. Partial [271, 167, 405, 507, 586, 596]. partially [11, 216, 417, 276, 434, 136]. parties [436]. partition [110]. Partitioning [290]. party [112]. Pascal [387]. Pascal-like [387]. passing [308]. path [273, 603, 126]. paths [428]. Patricia [603]. patterns [462, 231, 502]. Patterson [68]. PDL [482]. pebble [185]. pebbling [244]. perfect [476]. period [73]. Periodic [211, 73, 262]. Permutation [65, 368, 394]. permutations [146]. Petri [142, 114, 382, 163, 538, 205, 524, 92]. picture [330, 35]. pictures [618]. piecewise [555]. pigeonhole [500]. planar [352, 470, 614]. plane [479, 573]. point [407, 290, 204]. points [485, 515, 184, 622, 195]. poly [334, 377]. poly-time [377]. polymorphism [445]. polynomial [307, 387, 424, 363, 219, 39, 400, 314, 125, 43, 249, 95, 501, 623, 348, 72, 372]. polynomial-size [39, 501]. polynomial-time [400, 249, 623]. polynomials [424, 417, 235, 87, 150, 520, 325]. pomsets [480]. poset [261]. positive [129]. Possible [137, 547]. post [340]. post-conditions [340]. Power [499, 542, 416, 411, 67, 347, 597, 501, 180, 627]. power-free [67]. powerdomain [418]. powerdomains [33]. practice [637]. PRAM [625]. PRAMs [259]. pre [340]. pre- [340]. prefix [403]. Presburger [400]. presentations [225]. presented [638, 565]. primes [325]. primitive [325]. Principal [449, 411]. principles [500, 360]. priorities [601]. Probabilistic [342, 583, 407, 291, 151, 438, 64]. probability [548, 550, 510]. problem [38, 612, 76, 425, 163, 607, 21, 614, 135, 297, 397, 14, 193, 20, 110, 186, 353, 107, 289, 265, 518]. Problems [538, 504, 633, 371, 368, 219, 513, 178, 276, 281, 214, 117, 394, 330, 574, 248, 140, 229, 563, 15, 455, 310]. procedures [341, 231]. Process [79, 460, 515]. processes [86, 51, 571, 127, 285, 602, 196, 288, 343, 204, 437]. processing [532, 627]. processors [529]. product [424, 468, 269, 227]. products [256, 551]. program [53, 153, 581, 380]. Programming [339, 640, 613, 34, 271, 527, 281, 215, 628, 461, 455, 282, 636]. Programs [536, 481, 539, 387, 166, 91, 217, 447, 514, 621, 84, 118, 501, 345, 616, 438, 299, 459]. Projection [275]. Projections [540, 97]. PROLOG [359]. Proof [223, 138, 413, 162, 221, 179, 198, 22, 85]. proof-theoretic [85]. proofs [542, 481, 86, 500, 333]. Properties [259, 121, 407, 83, 18, 416, 552, 182, 279, 518]. property [387, 105, 7, 90, 143, 603, 484, 280]. Propositional [27, 612, 611, 567, 446]. protocols [102, 112, 361]. prover [9]. Proving [391, 359, 13]. pseudo [228, 582]. pseudo-c [582]. pseudorandom [257]. PSPACE [334]. PSPACE/ [334]. PT [521]. PT-nets [521]. public [564, 274]. public-key [564, 274]. pump [402]. pumping [209, 420, 28, 498]. pure [559]. purely [316]. push [270]. push-down [270]. Pushdown [155, 590, 50, 522].

Q [556]. Q-Adic [556]. quantitative [401]. queries [494]. query [119, 627]. question [403]. questions [94, 303]. queue [335]. queueing [534]. queues [426]. quotient [496].

races [471]. Random [168, 343, 128]. randomly [548]. rank [486]. rapid [243]. Rational [550, 230, 414, 107, 545, 104, 589, 427, 255, 276, 131, 420, 213, 347, 436, 519, 195]. rationality [584]. Reachability [205, 606, 540]. Real [228, 547, 28, 251, 379, 522]. Real-time [228, 547, 28, 522]. reasoning [161]. recognition [511, 88, 145, 466]. Recognizable [11, 131, 434]. recognizing [108]. reconnaissables [436]. reconnaissance [466]. Recurrent [1, 395]. recursion [450, 71, 239, 61]. recursion-theoretic [71]. Recursive [627, 499, 153, 341, 126, 494]. Recursively [360, 496, 16]. reduced [139]. reduces [73]. Reducibility [123]. reduction [137, 388, 148, 348, 381]. Reductions [52, 95, 309, 432]. Refusal [298]. register [190]. registers [157]. regular [264, 371, 217, 153, 79, 209, 584, 621, 75, 37, 484, 19, 69, 204, 469, 193, 366, 31, 324, 2, 175, 565, 212]. Regularity [262]. regulators [113]. rejection [534]. related [121, 281]. relatifs [549]. Relating [416]. relation [254]. Relational [166, 628, 364]. relations [113, 309, 420, 88, 230, 267, 512, 525, 72]. relationships [459]. relative [504, 549, 94]. relativized [620]. Relativizing [620]. remark [397, 626]. remarks [96]. Repeatable [356]. Repeated [286, 570]. Repetition [188]. Repetition-free [188]. repetitions [200]. replacement [52, 304]. representation [350, 496, 251, 379, 69, 17, 65, 519]. Representations [384, 479, 62, 577]. required [312]. residue [157]. Resolution [500, 206, 443, 541, 100, 81]. restricted [612]. restriction [596]. Restructuring [489]. result [498]. results [559, 430, 14]. retrieval [273]. Reversal [318, 378, 46]. Reversal-bounded [318, 46]. revisited [457]. rewrite [602, 472]. Rewriting [389, 598, 601, 111, 600, 383, 223, 323, 488, 597, 5, 229, 586, 280, 473, 415, 558]. right [270]. right-boundary [270]. ring [346, 361]. rings [157]. Robust [109, 322]. Rosser [623, 565, 105, 7, 90, 280, 278, 558]. rotation [579]. rotations [551]. routing [353]. rule [306, 601]. rules [602].

safe [388]. Sardinas [68]. Sardinas-Patterson [68]. Satisfiability [327, 612]. satisfy [101]. Satz [483]. SCCS [56]. scheduling [55, 201, 531]. Schematization [602]. scheme [10, 449]. schemes [491, 153, 342, 380]. Sci [634]. science [635, 82]. Scott [173]. search [425, 430, 574, 250, 231]. Searching [244, 510, 569]. second [581, 50, 212]. second-order [50, 212]. section [463]. security [102]. Selected [638]. selection [594]. selective [49, 99]. Self [245, 246]. Self-embedding [245, 246]. Semantic [161, 460, 254]. Semantics [215, 616, 206, 413, 86, 166, 197, 374, 610, 393, 390, 630, 560, 317, 283]. Semi [524, 128, 189, 246]. Semi-commutations [524]. semi-conditional [128]. semi-Thue [189, 246]. semicomputable [296]. semifunctors [134]. semigroup [312]. Semigroups [441, 209, 518]. semilattices [176]. semilinear [115, 338, 540]. semiring [121]. sensitive [115, 338, 191]. sentences [317]. sentential [327]. separating [625]. sequence [545, 518]. sequences [578, 547, 73, 41, 257, 556, 344, 345, 467]. Sequential [382, 266, 332, 631, 122, 25, 437]. serializability [60]. series [411, 180, 136]. server [335]. Set [490, 465, 371, 607, 470, 291, 487, 351]. set-covering [371]. set-union [607]. sets [165, 307, 384, 403, 290, 428, 4, 169, 599, 262, 242, 96, 602, 540, 267, 31, 568, 68, 317, 174, 495]. shift [190, 157]. shift-register [190]. shift-registers [157]. shortest [573, 410]. shuffle [315, 75]. shuffles [243]. Simon [70]. simple [604, 4, 220, 162, 242, 179, 42, 22]. Simulating [151]. simulation [471, 266]. simulations [604, 562]. Simultaneous [377, 1, 312]. single [335, 110]. single-server [335]. six [297]. six-state [297]. size [39, 592, 501, 236]. skew [551]. Skolem [162]. Skolem-Mahler-Lech [162]. SLD [443]. SLD-resolution [443]. slice [194, 63, 170]. slowly [15]. small [152]. smallest [108]. Smyth [418]. snapshots [286]. sofic [619, 557]. Software [82, 637]. solution [297, 110]. solutions [534, 187, 568, 233]. Solving [281, 87, 103, 615, 186]. Some [214, 96, 623, 94, 459, 239, 518, 612, 11, 224, 219, 216, 276, 48, 529, 556, 89, 455, 2, 303]. sort [221]. sorting [572, 179, 331]. sound [560]. space [604, 152, 377, 179, 247, 310]. space-bounded [604]. spanning [442]. sparse [530]. special [513, 90, 278, 325]. specifiability [97]. Specification [175, 517, 405, 629]. specifications [296, 167, 308, 20, 391, 154]. specified [595]. specifiés [595]. spectral [556]. Speeding [527]. Split [492]. Split-freedom [492]. squad [297]. square [284]. square-free [284]. st [250]. st-numbering [250]. Stable [503]. stacks [426]. stamp [102]. Stanley [30]. Star [174, 103]. Star-free [174]. state [631, 156, 386, 297, 438, 391]. state-independent [156]. states [214, 152]. Stationary [335]. Statistic [552]. Steiner [351]. step [395]. stochastic [120]. storage [98]. strategies [585]. streams [418, 283]. strict [522]. strictly [73]. string [320, 172, 597, 586]. string-rewriting [597, 586]. strings [462, 262]. strong [559]. Structural [512]. structure [475, 394, 273, 116, 260, 71, 415]. Structured [154, 34]. structures [446, 605, 123, 134, 117, 168, 343]. study [436]. subalgebras [36]. subclass [12, 469]. Subclasses [400]. subfamily [115, 338]. Subgraph [516, 76, 614]. subject [462]. Submodule [615]. submonoids [483]. subset [131]. subsets [11, 276, 434]. SUBST [223]. Substitution [457, 356, 99, 317]. substitutions [165, 549, 49, 274]. subsystems [559]. subtrees [548]. Subword [241, 554]. subwords [108]. Succinct [69]. sufficient [144]. suitability [86]. sum [312]. sums [572, 551]. superposition [9]. superstrings [410]. sups [173]. syllogisms [490]. symbolic [546]. symmetric [39, 235]. symposium [635]. synchronising [56]. synchronization [123, 297, 183]. synchronizing [464]. synchronous [346, 286, 287]. syntactic [104, 332, 631, 436]. syntax [301]. syntaxique [436]. Synthesis [433, 91, 217]. system [465, 203, 618, 223, 21, 540, 401, 376, 565, 558]. Systémes [549, 595]. systems [114, 601, 385, 111, 66, 181, 619, 286, 74, 600, 153, 222, 189, 52, 18, 262, 451, 214, 25, 171, 120, 87, 629, 7, 90, 411, 1, 433, 597, 581, 537, 193, 229, 623, 246, 586, 280, 595, 557, 93, 535, 293, 473, 567, 415, 472, 544, 549, 639]. systems-applications [153]. systolic [385, 54, 392].

T0L [1]. tables [101]. tape [116]. TAPSOFT [637]. Techniques [530, 201]. Technology [82]. Temporal [446, 542, 538, 77, 289, 254, 369, 421]. temps [466]. tensors [561]. Term [601, 91, 600, 280]. Term-rewriting [601, 600, 280]. termination [223, 438, 583]. terms [526, 608, 395, 261]. ternary [471]. tessellation [579]. Test [165]. Testing [573, 80, 349, 53, 7, 298]. tests [6]. text [108, 258, 466]. texte [466]. their [286, 416, 360, 409, 380]. theorem [539, 305, 9, 162, 359, 609, 483, 68, 558]. theorems [417, 70]. Theoret [634]. theoretic [71, 85]. Theoretical [82, 577]. theories [272, 169, 186, 146, 337, 301]. Theory [458, 62, 638, 413, 465, 382, 53, 199, 137, 370, 332, 631, 34, 480, 55, 507, 285, 608, 260, 364, 637, 493, 204, 50, 575, 596, 366, 49, 459, 398, 61]. thinning [158]. Third [82]. Thirteenth [636]. three [281, 143, 106]. three-dimensional [281]. three-element [143, 106]. Thue [189, 7, 90, 21, 193, 623, 246, 565, 518]. time [307, 547, 425, 604, 228, 219, 400, 281, 28, 377, 179, 247, 249, 13, 116, 431, 516, 297, 623, 110, 312, 376, 348, 466, 369, 522, 310, 72, 372]. time-bounded [604]. time-space [179, 247]. timed [437]. TMs [604]. tokens [142]. tool [54]. Topological [54, 329, 158, 595]. topologique [595]. topology [241]. TOPOS [523]. Total [299, 113]. tournament [487]. tours [573]. trace [86, 221, 453, 324]. traces [458]. tractable [19]. trade [179]. trade-off [179]. tradeoff [425]. Tradeoffs [511, 247]. Training [578]. transactions [60, 493]. transducer [155]. Transducers [200, 232, 522]. transductions [88, 506, 248, 17]. Transfer [417]. transform [319]. transformation [618, 221, 555]. transformations [153, 54, 462, 599]. Transforming [25]. transforms [588]. transition [222, 451, 535]. Transitive [121, 589, 439]. Translation [514]. transversals [4]. treatment [340]. tree [548, 463, 155, 350, 392, 52, 215, 329, 351, 149, 195]. Trees [339, 130, 44, 452, 205, 430, 603, 574, 151, 435, 373, 65, 177, 412, 64, 442]. triadic [561]. triangulation [574]. Tucker [591]. Turing [152, 592, 116, 164, 378]. Twelfth [282]. Two [406, 247, 213, 390, 630, 8, 98, 424, 132, 508, 112, 141, 496, 281, 73, 592, 579, 502, 243, 441, 522, 576, 472]. two- [281]. two-adjacent [141]. Two-dimensional [406, 592, 579]. Two-level [390, 630]. two-party [112]. Two-way [8, 98, 508, 579]. two-word [132]. Type [61, 206, 517, 137, 202, 295, 507, 164, 218, 20, 279, 449, 301]. type-0 [279]. typed [91, 624, 321]. types [296, 207, 203, 526, 218, 95, 300]. Typing [182].

unambiguous [309, 464, 236]. Unary [634, 240]. unbounded [450]. undecidability [246]. undecidable [454]. Unification [404, 599, 449, 399]. unified [506, 260, 49]. unifiers [169]. uniform [168, 283]. union [607]. unique [233]. Units [278]. universal [311, 152, 386]. universe [23]. unrepetitive [396]. unrestricted [19]. unsolvable [261]. Untermonoiden [483]. until [421]. update [425]. updating [442]. use [270]. using [359, 77]. utilized [15].

valued [232, 248, 327]. variable [203]. variables [570]. variant [128]. variants [224]. Varieties [177]. various [159, 581]. vector [214, 540, 358]. verifiability [112]. verification [481, 581, 77, 175]. version [560]. versions [612]. versus [583]. via [121]. viewpoint [603]. views [10]. VLSI [408]. vs [443].

walks [343]. way [508, 268, 579, 249, 8, 98]. Weak [581]. weaker [144]. Weakly [569]. weight [574]. weighted [435, 365]. Weyl [551]. where [73]. whether [131, 565]. which [115, 338, 235, 522]. While [118, 440]. While-programs [118]. width [501]. without [504, 429, 21, 386]. witness [96]. Word [321, 132, 21, 160, 258, 265]. words [633, 184, 403, 463, 70, 188, 1, 554, 356, 8, 436, 99, 243]. Worst [607]. Worst-case [607]. WRAMs [259]. writes [312].

X [594].

Yacc [326]. years [203].

zeta [584].

References

Lange:1985:RWS

[1] K.-J. Lange and E. Welzl. Recurrent words and simultaneous growth in T0L systems. Theoretical Computer Science, 35(1):1–15, January 1985. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stern:1985:CSC

[2] J. Stern. Characterizations of some classes of regular events. Theoretical Computer Science, 35(1):17–42, January 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bloom:1985:LCO

[3] S. L. Bloom and D. R. Troeger. A logical characterization of observation equivalence. Theoretical Computer Science, 35 (1):43–53, January 1985. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Edelsbrunner:1985:FTS

[4] H. Edelsbrunner. Finding transversals for sets of simple geometric figures. Theoretical Computer Science, 35(1):55– 69, January 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Metivier:1985:CLR

[5] Y. Metivier. Calculation of the length of rewriting chains in free monoide. Theoretical Computer Science, 35(1):71– 87, January 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Benoit:1985:AT

[6] C. Benoit. Axiomatisation of tests. Theoretical Computer Science, 35(1):89– 107, January 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kapur:1985:ATC

[7] D. Kapur, M. S. Krishnamoorthy, R. McNaughton, Narendran, and P. An O(modT mod ^{3}) algorithm for testing the Church–Rosser property of Thue systems. Theoretical Computer Science, 35(1):109–114, January 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pecuchet:1985:TWF

[8] J.-P. Pecuchet. Two-way finite automata and finite words. Theoretical Computer Science, 35(1):115–122, January 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fribourg:1985:SOT

[9] L. Fribourg. A superposition oriented theorem prover. Theoretical Computer Science, 35(2–3):129–164, February 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ausiello:1985:EAV

[10] G. Ausiello, A. d’Atri, and M. Moscarini. On the existence of acyclic views in a database scheme. Theoretical Computer Science, 35(2–3):165–177, February 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cori:1985:RSS

[11] R. Cori and Y. Metivier. Recognizable subsets of some partially Abelian monoids. Theoretical Computer Science, 35(2–3):179–189, February 1985. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Memmi:1985:IFN

[12] G. Memmi and A. Finkel. An introduction to FIFO nets-monogeneous nets: a subclass of FIFO nets. Theoretical Computer Science, 35(2–3):191–214, February 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kobayashi:1985:PTC

[13] K. Kobayashi. On proving time constructibility of functions. Theoretical Computer Science, 35(2–3):215–225, February 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Narendran:1985:CRC

[14] P. Narendran and F. Otto. Complexity results on the conjugacy problem for monoids. Theoretical Computer Science, 35(2–3):227–243, February 1985. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Plaisted:1985:CDP

[15] D. A. Plaisted. Complete divisibility problems for slowly utilized oracles. Theoretical Computer Science, 35(2–3): 245–260, February 1985. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hirose:1985:HCR

[16] S. Hirose, S. Okawa, and M. Yoneda. A homomorphic characterization of recursively enumerable languages. Theoretical Computer Science, 35(2–3):261– 269, February 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pin:1985:AMR

[17] J.-E. Pin and J. Sakarovitch. An application of the matrix representation of transductions. Theoretical Computer Science, 35(2–3):271–293, February 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Head:1985:CPH

[18] T. Head and J. Wilkinson. Code properties and homomorphisms of D0L systems. Theoretical Computer Science, 35 (2–3):295–312, February 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Leiss:1985:CTU

[19] E. Leiss. On classes of tractable unrestricted regular expressions. Theoretical Computer Science, 35(2–3):313– 327, February 1985. CODEN TCSCDI.

ISSN 0304-3975 (print), 1879-2294 (electronic).

Oyamaguchi:1985:DTE

[20] M. Oyamaguchi. On the data type extension problem for algebraic specifications. Theoretical Computer Science, 35 (2–3):329–336, February 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kapur:1985:FTS

[21] D. Kapur and P. Narendran. A finite Thue system with decidable word problem and without equivalent finite canonical system. Theoretical Computer Science, 35(2–3):337–344, February 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sain:1985:SPC

[22] I. Sain. A simple proof for the completeness of Floyd’s method. Theoretical Computer Science, 35(2–3):345– 348, February 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Broy:1985:HKU

[23] M. Broy. On the Herbrand-Kleene universe for nondeterministic computations. Theoretical Computer Science, 36 (1):1–19, March 1985. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Engelfriet:1985:DOE

[24] J. Engelfriet. Determinacy to (observation equivalence=trace equivalence). Theoretical Computer Science, 36(1):21– 25, March 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Janicki:1985:TSS

[25] R. Janicki. Transforming sequential systems into concurrent systems. Theoretical Computer Science, 36(1):27– 58, March 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blum:1985:LBM

[26] N. Blum. An Ω(n^{4}/3) lower bound on the monotone network complexity of the n^{th} degree convolution. Theoretical Computer Science, 36(1):59–69, March 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tiomkin:1985:PDL

[27] M. L. Tiomkin and J. A. Makowsky. Propositional dynamic logic with local assignments. Theoretical Computer Science, 36(1):71–87, March 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Igarashi:1985:PLR

[28] Y. Igarashi. A pumping lemma for real-time deterministic context-free languages. Theoretical Computer Science, 36(1):89–97, March 1985. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DeFelice:1985:CFC

[29] C. De Felice. Construction of factorisation codes. Theoretical Computer Science, 36(1):99–108, March 1985. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hirose:1985:CSH

[30] S. Hirose and M. Yoneda. On the Chomsky and Stanley’s homomorphic characterization of context-free languages. Theoretical Computer Science, 36(1): 109–112, March 1985. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ruohonen:1985:EMS

[31] K. Ruohonen. On equality of multiplicity sets of regular languages. Theoretical Computer Science, 36(1):113– 117, March 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Scarpellini:1985:CBN

[32] B. Scarpellini. Complex Boolean networks obtained by diagonalization. Theoretical Computer Science, 36(1):119– 125, March 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Winskel:1985:PM

[33] G. Winskel. On powerdomains and modality. Theoretical Computer Science, 36(1):127–137, March 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fenton:1985:GMT

[34] N. E. Fenton, R. W. Whitty, and A. A. Kaposi. A generalised mathematical theory of structured programming. Theoretical Computer Science, 36(2–3):145– 171, April 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sudborough:1985:CDC

[35] I. H. Sudborough and E. Welzl. Complexity and decidability for chain code picture languages. Theoretical Computer Science, 36(2–3):173–202, April 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Demel:1985:FAC

[36] J. Demel, M. Demlova, and V. Koubek. Fast algorithms constructing minimal subalgebras, congruences, and ideals in a finite algebra. Theoretical Computer Science, 36(2–3):203–216, April 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kaminski:1985:COR

[37] M. Kaminski. A classification of omegaregular languages. Theoretical Computer Science, 36(2–3):217–229, April 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Allender:1985:ILB

[38] E. Allender and M. M. Klawe. Improved lower bounds for the cycle detection problem. Theoretical Computer Science, 36(2–3):231–237, April 1985. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fagin:1985:BDP

[39] R. Fagin, M. M. Klawe, N. J. Pippenger, and L. Stockmeyer. Bounded-depth, polynomial-size circuits for symmetric functions. Theoretical Computer Science, 36(2–3):239–250, April 1985. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

FarinasDelCerro:1985:DLD

[40] L. Farinas Del Cerro and E. Orlowska. DAL-a logic for data analysis. Theoretical Computer Science, 36(2–3):251– 264, April 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jerrum:1985:CFM

[41] M. R. Jerrum. The complexity of finding minimum-length generator sequences. Theoretical Computer Science, 36(2–3):265–289, April 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Matsuno:1985:ASM

[42] H. Matsuno, K. Inoue, H. Taniguchi, and I. Takanami. Alternating simple multihead finite automata. Theoretical Computer Science, 36(2–3):291– 308, April 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Keller-Gehrig:1985:FAC

[43] W. Keller-Gehrig. Fast algorithms for the characteristic polynomial. Theoretical Computer Science, 36(2–3):309– 317, April 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diks:1985:EBT

[44] K. Diks. Embeddings of binary trees in lines. Theoretical Computer Science, 36(2–3):319–331, April 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alt:1985:MEN

[45] H. Alt. Multiplication is the easiest nontrivial arithmetic function. Theoretical Computer Science, 36(2–3):333– 339, April 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rytter:1985:CRB

[46] W. Rytter and M. Chrobak. A characterization of reversal-bounded multipushdown machine languages. Theoretical Computer Science, 36(2–3):341– 344, April 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Farr:1985:CCH

[47] G. Farr and C. McDiarmid. The complexity of counting homeomorphs. Theoretical Computer Science, 36(2–3):345– 348, April 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ko:1985:SNC

[48] K. Ko. On some natural complete operators. Theoretical Computer Science, 37 (1):1–30, May 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:1985:CSS

[49] G. Rozenberg. On coordinated selective substitutions: towards a unified theory of grammars and machines. Theoretical Computer Science, 37(1):31–50, May 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Muller:1985:TEP

[50] D. E. Muller and P. E. Schupp. The theory of ends, pushdown automata, and second-order logic. Theoretical Computer Science, 37(1):51–75, May 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bergstra:1985:ACP

[51] J. A. Bergstra and J. W. Klop. Algebra of communicating processes with abstraction. Theoretical Computer Science, 37(1):77–121, May 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gallier:1985:RTR

[52] J. H. Gallier and R. V. Book. Reductions in tree replacement systems. Theoretical Computer Science, 37(2):123– 150, November 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bouge:1985:CTP

[53] L. Bouge. A contribution to the theory of program testing. Theoretical Computer Science, 37(2):151–181, November 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Culik:1985:TTT

[54] K. Culik, II and I. Fris. Topological transformations as a tool in the design of systolic networks. Theoretical Computer Science, 37(2):183–216, November 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gonczarowski:1985:AST

[55] J. Gonczarowski and M. K. Warmuth. Applications of scheduling theory to formal language theory. Theoretical Computer Science, 37(2):217–243, November 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deSimone:1985:HLS

[56] R. de Simone. Higher-level synchronising devices in MEIJE-SCCS. Theoretical Computer Science, 37(3):245– 267, December 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tarlecki:1985:EFM

[57] A. Tarlecki. On the existence of free models in abstract algebraic institutions. Theoretical Computer Science, 37 (3):269–304, December 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Darondeau:1985:AFA

[58] P. Darondeau. About fair asynchrony. Theoretical Computer Science, 37(3): 305–336, December 1985. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ehrenfeucht:1985:AGF

[59] A. Ehrenfeucht, H. C. M. Kleijn, and G. Rozenberg. Adding global forbidding context to context-free grammars. Theoretical Computer Science, 37(3):357– 360, December 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fle:1985:MSI

[60] M. P. Fle and G. Roucairol. Maximal serializability of iterated transactions. Theoretical Computer Science, 38 (1):1–16, May 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Weihrauch:1985:TRT

[61] K. Weihrauch. Type 2 recursion theory. Theoretical Computer Science, 38 (1):17–33, May 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kreitz:1985:TR

[62] C. Kreitz and K. Weihrauch. Theory of representations. Theoretical Computer Science, 38(1):35–53, May 1985. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wegener:1985:CSF

[63] I. Wegener. On the complexity of slice functions. Theoretical Computer Science, 38(1):55–68, May 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Snir:1985:LBP

[64] M. Snir. Lower bounds on probabilistic linear decision trees. Theoretical Computer Science, 38(1):69–82, May 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ramanan:1985:PRA

[65] P. V. Ramanan and C. L. Liu. Permutation representation of k-ary trees. Theoretical Computer Science, 38(1): 83–98, May 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beeri:1985:FSJ

[66] C. Beeri and M. Y. Vardi. Formal systems for join dependencies. Theoretical Computer Science, 38(1):99–116, May 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Leconte:1985:CPF

[67] M. Leconte. A characterization of power-free morphisms. Theoretical Computer Science, 38(1):117–122, May 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Van:1985:CCS

[68] D. L. Van. Code-compatible sets and a generalisation of the Sardinas-Patterson theorem. Theoretical Computer Science, 38(1):123–132, May 1985. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Leiss:1985:SRR

[69] E. Leiss. Succinct representation of regular languages by Boolean automata. II. Theoretical Computer Science, 38(1): 133–136, May 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Finkel:1985:GTH

[70] A. Finkel. A generalisation of the theorems of Higman and of Simon to infinite words. Theoretical Computer Science, 38(1):137–142, May 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schmidt:1985:RTS

[71] D. Schmidt. The recursion-theoretic structure of complexity classes. Theoretical Computer Science, 38(2–3):143– 156, June 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Watanabe:1985:OOP

[72] O. Watanabe. On one-one polynomial time equivalence relations. Theoretical Computer Science, 38(2–3):157– 165, June 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Imori:1985:PCS

[73] M. Imori and H. Yamada. Periodic character sequences where identifying two characters strictly reduces the period. Theoretical Computer Science, 38(2–3): 167–192, June 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Burkhard:1985:ICC

[74] H. D. Burkhard. An investigation of controls for concurrent systems based on abstract control languages. Theoretical Computer Science, 38(2–3):193– 122, June 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jantzen:1985:ERE

[75] M. Jantzen. Extending regular expressions with iterated shuffle. Theoretical Computer Science, 38(2–3):223– 247, June 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Asano:1985:ASH

[76] T. Asano. An approach to the subgraph homeomorphism problem. Theoretical Computer Science, 38(2–3):249– 267, June 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mishra:1985:HVA

[77] B. Mishra and E. Clarke. Hierarchical verification of asynchronous circuits using temporal logic. Theoretical Computer Science, 38(2–3):269–291, June 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gonzalez:1985:CMM

[78] T. F. Gonzalez. Clustering to minimize the maximum intercluster distance. Theoretical Computer Science, 38(2–3): 293–306, June 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Harel:1985:PLR

[79] D. Harel and D. Peleg. Process logic with regular formulas. Theoretical Computer Science, 38(2–3):307–322, June 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rosenkrantz:1985:TGC

[80] D. J. Rosenkrantz and H. B. Hunt, III. Testing for grammatical coverings. Theoretical Computer Science, 38(2–3):323– 341, June 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Linial:1985:DHC

[81] N. Linial and M. Tarsi. Deciding hypergraph 2-colourability by H-resolution. Theoretical Computer Science, 38(2–3): 343–347, June 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:1985:TCF

[82] Anonymous. Third Conference on Foundations of Software Technology and Theoretical Computer Science. Theoretical Computer Science, 39(1):??, July 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Frougny:1985:CFG

[83] C. Frougny. Context-free grammars with cancellation properties. Theoretical Computer Science, 39(1):3–13, July 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lassez:1985:OFL

[84] J.-L. Lassez and M. J. Maher. Optimal fixedpoints of logic programs. Theoretical Computer Science, 39(1):15–25, July 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stirling:1985:PTC

[85] C. Stirling. A proof-theoretic characterization of observational equivalence. Theoretical Computer Science, 39(1): 27–45, July 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Back:1985:STS

[86] R. J. R. Back and H. Mannila. On the suitability of trace semantics for modular proofs of communicating processes. Theoretical Computer Science, 39 (1):47–68, July 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kannan:1985:SSL

[87] R. Kannan. Solving systems of linear equations over polynomials. Theoretical Computer Science, 39(1):69–88, July 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ibarra:1985:ERT

[88] O. H. Ibarra, M. A. Palis, and J. H. Chang. On efficient recognition of transductions and relations. Theoretical Computer Science, 39(2–3):89– 106, August 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Priese:1985:FDM

[89] L. Priese. On a fast decomposition method in some models of concurrent computations. Theoretical Computer Science, 39(2–3):107–121, August 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kapur:1985:CRP

[90] D. Kapur, P. Narendran, M. S. Krishnamoorthy, McNaughton, and R. The Church–Rosser property and special Thue systems. Theoretical Computer Science, 39(2–3):123–133, August 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bohm:1985:AST

[91] Corrado Böhm and Alessandro Berarducci. Automatic synthesis of typed Λ-programs on term algebras. Theoretical Computer Science, 39(2–3):135– 153, August 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Parigot:1985:LAP

[92] M. Parigot and E. Pelz. A logical approach of Petri net languages. Theoretical Computer Science, 39(2–3):155– 169, August 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Spehner:1985:ESE

[93] J.-C. Spehner. Entire systems of equations on a finite alphabet and Ehrenfeucht’s conjecture. Theoretical Computer Science, 39(2–3):171–188, August 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

RodriguezArtalejo:1985:SQA

[94] M. Rodriguez Artalejo. Some questions about expressiveness and relative completeness in Hoare’s logic. Theoretical Computer Science, 39(2–3):189– 206, August 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mahaney:1985:RAP

[95] S. R. Mahaney and P. Young. Reductions among polynomial isomorphism types. Theoretical Computer Science, 39 (2–3):207–224, August 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Joseph:1985:SRW

[96] D. Joseph and P. Young. Some remarks on witness functions for nonpolynomial and noncomplete sets in NP. Theoretical Computer Science, 39(2–3):225– 237, August 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hull:1985:NFS

[97] R. Hull. Non-finite specifiability of projections of functional dependency families. Theoretical Computer Science, 39 (2–3):239–265, August 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vogel:1985:TWA

[98] J. Vogel and K. Wagner. Two-way automata with more than one storage medium. Theoretical Computer Science, 39(2–3):267–280, August 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Siromoney:1985:IWO

[99] R. Siromoney and V. R. Dare. On infinite words obtained by selective substitution grammars. Theoretical Computer Science, 39(2–3):281–295, August 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Haken:1985:IR

[100] A. Haken. The intractability of resolution. Theoretical Computer Science, 39 (2–3):297–308, August 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ginsburg:1985:CTS

[101] S. Ginsburg and E. H. Spanier. On completing tables to satisfy functional dependencies. Theoretical Computer Science, 39(2–3):309–317, August 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Book:1985:SNS

[102] R. V. Book and F. Otto. On the security of name-stamp protocols. Theoretical Computer Science, 39(2–3):319– 325, August 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Leiss:1985:SSE

[103] E. L. Leiss. On solving star equations. Theoretical Computer Science, 39(2–3): 327–332, August 1985. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Arnold:1985:SCR

[104] A. Arnold. A syntactic congruence for rational omega-languages. Theoretical Computer Science, 39(2–3):333– 335, August 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bunder:1985:EKC

[105] M. W. Bunder. An extension of Klop’s counterexample to the Church– Rosser property to lambda-calculus with other ordered pair combinators. Theoretical Computer Science, 39(2–3):337– 342, August 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Karhumaki:1985:TEC

[106] J. Karhumaki. On three-element codes.

Theoretical Computer Science, 40(1):3– 11, September 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Restivo:1985:RLB

[107] A. Restivo and C. Reutenauer. Rational languages and the Burnside problem. Theoretical Computer Science, 40(1):13– 30, September 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blumer:1985:SAR

[108] A. Blumer, J. Blumer, D. Haussler, A. Ehrenfeucht, Chen, M. T., and J. Seiferas. The smallest automation recognizing the subwords of a text. Theoretical Computer Science, 40(1):31–55, September 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schoning:1985:RAD

[109] U. Schoning. Robust algorithms: a different approach to oracles. Theoretical Computer Science, 40(1):57–66, September 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Paige:1985:LTS

[110] R. Paige, R. E. Tarjan, and R. Bonic. A linear time solution to the single function coarsest partition problem. Theoretical Computer Science, 40(1):67–84, September 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bauer:1985:LRS

[111] G. Bauer. n-level rewriting systems. Theoretical Computer Science, 40(2–3): 85–99, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Book:1985:VTA

[112] R. V. Book and F. Otto. On the verifiability of two-party algebraic protocols. Theoretical Computer Science, 40(2–3): 101–130, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bucher:1985:TRG

[113] W. Bucher, A. Ehrenfeucht, and D. Haussler. On total regulators generated by derivation relations. Theoretical Computer Science, 40(2–3):131– 148, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aalbersberg:1985:CSP

[114] I. J. J. Aalbersberg and G. Rozenberg. CTS systems and Petri nets. Theoretical Computer Science, 40(2–3):149– 162, ???? 1985. CODEN TCSCDI.

ISSN 0304-3975 (print), 1879-2294 (electronic).

Ayers:1985:DAS

[115] K. Ayers. Deque automata and a subfamily of context-sensitive languages which contains all semilinear bounded languages. Theoretical Computer Science, 40(2–3):163–174, ???? 1985. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kobayashi:1985:SOT

[116] K. Kobayashi. On the structure of onetape nondeterministic Turing machine time hierarchy. Theoretical Computer Science, 40(2–3):175–193, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Huang:1985:IFS

[117] Ming-Deh A. Huang and K. J. Lieberherr. Implications of forbidden structures for extremal algorithmic problems. Theoretical Computer Science, 40(2–3): 195–210, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mascari:1985:WPN

[118] G. Mascari and M. Venturini Zilli.

While-programs with nondeterministic assignments and the logic ALNA. Theoretical Computer Science, 40(2–3):211– 235, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Balcazar:1985:BQM

[119] J. L. Balcazar, R. V. Book, and U. Schoning. On bounded query machines. Theoretical Computer Science, 40(2–3):237–243, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kanaoka:1985:HDS

[120] T. Kanaoka and S. Tomita. Homogeneous decomposition of stochastic systems. Theoretical Computer Science, 40 (2–3):245–255, ???? 1985. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abdali:1985:TCR

[121] S. K. Abdali and B. D. Saunders. Transitive closure and related semiring properties via eliminants. Theoretical Computer Science, 40(2–3):257–274, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fogelman-Soulie:1985:PSC

[122] F. Fogelman-Soulie. Parallel and sequential computation on Boolean networks. Theoretical Computer Science, 40(2–3):275–300, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ginzburg:1985:RSS

[123] A. Ginzburg and M. Yoeli. Reducibility of synchronization structures. Theoretical Computer Science, 40(2–3):301– 314, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Urbanek:1985:GNF

[124] F. J. Urbanek. On Greibach normal form construction. Theoretical Computer Science, 40(2–3):315–317, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kaminski:1985:LBP

[125] M. Kaminski. A lower bound for polynomial multiplication. Theoretical Computer Science, 40(2–3):319–322, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Krishnamoorthy:1985:RPO

[126] M. S. Krishnamoorthy and P. Narendran. On recursive path ordering. Theoretical Computer Science, 40(2–3):323– 328, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gonthier:1985:ACP

[127] G. Gonthier. Algebraic calculi of processes and net expressions. Theoretical Computer Science, 40(2–3):329– 337, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Paun:1985:VRC

[128] G. Păun. A variant of random context grammars: semi-conditional grammars. Theoretical Computer Science, 41 (1):1–17, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Goles:1985:DPA

[129] E. Goles. Dynamics of positive automata networks. Theoretical Computer Science, 41(1):19–32, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cazanescu:1985:CFT

[130] V. E. Cazanescu. On context-free trees.

Theoretical Computer Science, 41(1):33– 50, ???? 1985. CODEN TCSCDI.

ISSN 0304-3975 (print), 1879-2294 (electronic).

Gohon:1985:ADW

[131] P. Gohon. An algorithm to decide whether a rational subset of N^{k} is recognizable. Theoretical Computer Science, 41(1):51–59, ???? 1985. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barbin-LeRest:1985:CTW

[132] E. Barbin-Le Rest and M. Le Rest. On the combinatorial of two-word codes. Theoretical Computer Science, 41(1):61– 80, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Iliopoulos:1985:CGA

[133] C. S. Iliopoulos. Computing in general Abelian groups is hard. Theoretical Computer Science, 41(1):81–93, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hayashi:1985:ASC

[134] S. Hayashi. Adjunction of semifunctors: categorical structures in nonextensional lambda calculus. Theoretical Computer Science, 41(1):95–104, ???? 1985. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Maon:1985:EPC

[135] Y. Maon. On the equivalence problem of compositions of morphisms and inverse morphisms on context-free languages. Theoretical Computer Science, 41(1):105–107, ???? 1985. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Thibon:1985:IAF

[136] J.-Y. Thibon. Integrity of algebras of formal series on a partially commutative alphabet. Theoretical Computer Science, 41(1):109–112, ???? 1985. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bunder:1985:PFE

[137] M. W. Bunder. Possible forms of evaluation or reduction in Martin-Löf type theory. Theoretical Computer Science, 41(1):113–120, ???? 1985. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Albert:1985:PEC

[138] M. H. Albert and J. Lawrence. A proof of Ehrenfeucht’s conjecture. Theoretical Computer Science, 41(1):121–123, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Helfrich:1985:ACM

[139] B. Helfrich. Algorithms to construct Minkowski reduced and Hermite reduced lattice bases. Theoretical Computer Science, 41(2–3):125–139, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Monien:1985:BCN

[140] B. Monien and I. H. Sudborough. Bandwidth constrained NP-complete problems. Theoretical Computer Science, 41 (2–3):141–167, ???? 1985. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dahlhaus:1985:CTA

[141] E. Dahlhaus and H. Gaifman. Concerning two-adjacent context-free languages.

Theoretical Computer Science, 41(2–3): 169–184, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Reisig:1985:PNI

[142] W. Reisig. Petri nets with individual tokens. Theoretical Computer Science, 41(2–3):185–213, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Karhumaki:1985:PTE

[143] J. Karhumaki. A property of threeelement codes. Theoretical Computer Science, 41(2–3):215–222, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tomita:1985:WSC

[144] E. Tomita and K. Seino. A weaker sufficient condition for the equivalence of a pair of DPDAs to be decidable. Theoretical Computer Science, 41(2–3):223– 230, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ibarra:1985:FPL

[145] O. H. Ibarra, M. A. Palis, and S. M. Kim. Fast parallel language recognition by cellular automata. Theoretical Computer Science, 41(2–3):231–246, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pelz:1985:CTP

[146] E. Pelz. On the complexity of theories of permutations. Theoretical Computer Science, 41(2–3):247–269, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Grant:1985:IND

[147] J. Grant and J. Minker. Inferences for numerical dependencies. Theoretical Computer Science, 41(2–3):271– 287, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hirokawa:1985:CCR

[148] S. Hirokawa. Complexity of the combinator reduction machine. Theoretical Computer Science, 41(2–3):289– 303, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Slutzki:1985:ATA

[149] G. Slutzki. Alternating tree automata.

Theoretical Computer Science, 41(2–3): 305–318, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stoss:1985:CEI

[150] H.-J. Stoss. The complexity of evaluating interpolation polynomials. Theoretical Computer Science, 41(2–3):319– 323, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

MeyerAufDerHeide:1985:SPD

[151] F. Meyer Auf Der Heide. Simulating probabilistic by deterministic algebraic computation trees. Theoretical Computer Science, 41(2–3):325–330, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Inoue:1985:ALT

[152] K. Inoue, I. Takanami, and R. Vollmar. Alternating on-line Turing machines with only universal states and small space bounds. Theoretical Computer Science, 41(2–3):331–339, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Courcelle:1986:ETR

[153] B. Courcelle. Equivalences and transformations of regular systems-applications to recursive program schemes and grammars. Theoretical Computer Science, 42 (1):1–122, May 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wirsing:1986:SAS

[154] M. Wirsing. Structured algebraic specifications: a kernel language. Theoretical Computer Science, 42(2):123– 249, August 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Engelfriet:1986:PMM

[155] J. Engelfriet and H. Vogler. Pushdown machines for the macro tree transducer. Theoretical Computer Science, 42(3):251–368, October 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Grzymala-Busse:1986:CSI

[156] J. W. Grzymala-Busse and Z. Bavel. Characterization of state-independent automata. Theoretical Computer Science, 43(1):1–10, May 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Scheuing:1986:DLA

[157] A. Scheuing. Decomposition of linear automata over residue rings into shiftregisters. Theoretical Computer Science, 43(1):11–30, May 1986. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ronse:1986:TCT

[158] C. Ronse. A topological characterization of thinning. Theoretical Computer Science, 43(1):31–41, May 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Goldschlager:1986:CPC

[159] L. M. Goldschlager and I. Parberry.

On the construction of parallel computers from various bases of Boolean functions. Theoretical Computer Science, 43 (1):43–58, May 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Redziejowski:1986:IWL

[160] R. R. Redziejowski. Infinite-word languages and continuous mappings. Theoretical Computer Science, 43(1):59– 79, May 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Orlowska:1986:SAI

[161] E. Orlowska. Semantic analysis of inductive reasoning. Theoretical Computer Science, 43(1):81–89, May 1986. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hansel:1986:DSP

[162] G. Hansel. The demonstration of a simple proof of skolem-mahler-lech theorem. Theoretical Computer Science, 43 (1):91–98, May 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Clote:1986:FCP

[163] P. Clote. On the finite containment problem for Petri nets. Theoretical Computer Science, 43(1):99–105, May 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lindsay:1986:AOT

[164] P. A. Lindsay. Alternation and omegatype Turing acceptors. Theoretical Computer Science, 43(1):107–115, May 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Albert:1986:TSF

[165] M. H. Albert and J. Lawrence. Test sets for finite substitutions. Theoretical Computer Science, 43(1):117–122, May 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berghammer:1986:RAS

[166] R. Berghammer and H. Zierer. Relational algebraic semantics of deterministic and nondeterministic programs. Theoretical Computer Science, 43(2–3):123– 147, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Heering:1986:PEO

[167] J. Heering. Partial evaluation and omega-completeness of algebraic specifications. Theoretical Computer Science, 43(2–3):149–167, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jerrum:1986:RGC

[168] M. R. Jerrum, L. G. Valiant, and V. V. Vazirani. Random generation of combinatorial structures from a uniform distribution. Theoretical Computer Science, 43(2–3):169–188, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fages:1986:CSU

[169] F. Fages and G. Huet. Complete sets of unifiers and matchers in equational theories. Theoretical Computer Science, 43(2–3):189–200, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wegener:1986:MCS

[170] I. Wegener. More on the complexity of slice functions. Theoretical Computer Science, 43(2–3):201–211, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Janicki:1986:CMC

[171] R. Janicki, P. E. Lauer, M. Koutny, and R. Devillers. Concurrent and maximally concurrent evolution of nonsequential systems. Theoretical Computer Science, 43(2–3):213–238, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Landau:1986:ESM

[172] G. M. Landau and U. Vishkin. Efficient string matching with k mismatches. Theoretical Computer Science, 43(2–3): 239–249, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pasztor:1986:SIC

[173] A. Pasztor and R. Statman. Scott induction and closure under omega-sups. Theoretical Computer Science, 43(2–3): 251–263, ???? 1986. CODEN TCSCDI.

ISSN 0304-3975 (print), 1879-2294 (electronic).

deLuca:1986:SFS

[174] A. de Luca and A. Restivo. Star-free sets of integers. Theoretical Computer Science, 43(2–3):265–275, ???? 1986. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Suzuki:1986:SVD

[175] I. Suzuki, Y. Motohashi, K. Taniguchi, T. Kasami, and T. Okamoto. Specification and verification of decentralized daisy chain arbiters with omegaextended regular expressions. Theoretical Computer Science, 43(2–3):277– 291, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Adamek:1986:CS

[176] J. Adamek, J. Reiterman, and E. Nelson. Continuous semilattices. Theoretical Computer Science, 43(2–3):293– 313, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Saoudi:1986:VAD

[177] A. Saoudi. Varieties of automata deriving from infinite trees. Theoretical Computer Science, 43(2–3):315–335, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Edwards:1986:CCP

[178] K. Edwards. The complexity of colouring problems on dense graphs. Theoretical Computer Science, 43(2–3):337– 343, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Johnson:1986:SPT

[179] D. B. Johnson. A simple proof of a timespace trade-off for sorting with linear comparisons. Theoretical Computer Science, 43(2–3):345–350, ???? 1986. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ritzmann:1986:FNA

[180] P. Ritzmann. A fast numerical algorithm for the composition of power series with complex coefficients. Theoretical Computer Science, 44(1):1–16, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blanchard:1986:CS

[181] F. Blanchard and G. Hansel. Coded systems. Theoretical Computer Science, 44 (1):17–49, ???? 1986. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Leivant:1986:TCP

[182] D. Leivant. Typing and computational properties of lambda expressions. Theoretical Computer Science, 44(1):51– 68, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rosier:1986:BEC

[183] L. E. Rosier and H.-C. Yen. Boundedness, empty channel detection, and synchronization for communicating finite automata. Theoretical Computer Science, 44(1):69–105, ???? 1986. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blanchard:1986:DOP

[184] F. Blanchard and S. Martinez. Dense orbit points of certain languages of infinite words. Theoretical Computer Science, 44(1):107–110, ???? 1986. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chang:1986:PA

[185] J. H. Chang, O. H. Ibarra, M. A. Palis, and B. Ravikumar. On pebble automata. Theoretical Computer Science, 44(1):111–121, ???? 1986. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Paul:1986:SEP

[186] E. Paul. On solving the equality problem in theories defined by Horn clauses. Theoretical Computer Science, 44(2):127– 153, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Leiss:1986:GLE

[187] E. L. Leiss. Generalized language equations with multiple solutions. Theoretical Computer Science, 44(2):155– 174, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kobayashi:1986:RFW

[188] Y. Kobayashi. Repetition-free words. Theoretical Computer Science, 44(2): 175–197, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diekert:1986:CST

[189] V. Diekert. Complete semi-Thue systems for Abelian groups. Theoretical Computer Science, 44(2):199–208, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Etzion:1986:AGS

[190] T. Etzion. An algorithm for generating shift-register cycles. Theoretical Computer Science, 44(2):209–224, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Okawa:1986:IHC

[191] S. Okawa, S. Hirose, and M. Yoneda. On the impossibility of the homomorphic characterization of context-sensitive languages. Theoretical Computer Science, 44(2):225–228, ???? 1986. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Urbanek:1986:GNF

[192] F. J. Urbanek. On Greibach normal form construction. Theoretical Computer Science, 44(2):229–236, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Narendran:1986:EPR

[193] P. Narendran. On the equivalence problem for regular Thue systems. Theoretical Computer Science, 44(2):237– 245, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dunne:1986:CCS

[194] P. E. Dunne. The complexity of central slice functions. Theoretical Computer Science, 44(3):247–257, ???? 1986. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Takahashi:1986:GFP

[195] M. Takahashi. The greatest fixed-points and rational omega-tree languages. Theoretical Computer Science, 44(3):259– 274, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kreowski:1986:NPC

[196] H.-J. Kreowski and A. Wilharm. Net processes correspond to derivation processes in graph grammars. Theoretical Computer Science, 44(3):275–305, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dowling:1986:CSF

[197] W. F. Dowling and J. H. Gallier. Continuation semantics for flowgraph equations. Theoretical Computer Science, 44 (3):307–331, ???? 1986. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Poulsen:1986:ECA

[198] E. T. Poulsen. The Ehrenfeucht Conjecture: an algebra-framework for its proof. Theoretical Computer Science, 44 (3):333–339, ???? 1986. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Broy:1986:TNP

[199] M. Broy. A theory for nondeterminism, parallelism, communication, and concurrency. Theoretical Computer Science, 45(1):1–61, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Crochemore:1986:TR

[200] M. Crochemore. Transducers and repetitions. Theoretical Computer Science, 45(1):63–86, ???? 1986. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gonczarowski:1986:MDF

[201] J. Gonczarowski. Manipulating derivation forests by scheduling techniques. Theoretical Computer Science, 45(1):87– 119, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dezani-Ciancaglini:1986:CFC

[202] M. Dezani-Ciancaglini and I. Margaria.

A characterization of F-complete type assignments. Theoretical Computer Science, 45(2):121–157, ???? 1986. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Girard:1986:SFV

[203] J.-Y. Girard. The system F of variable types, fifteen years later. Theoretical Computer Science, 45(2):159–192, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Meyer:1986:MRP

[204] J.-J. Ch. Meyer. Merging regular processes by means of fixed-point theory. Theoretical Computer Science, 45(2): 193–260, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Huber:1986:RTH

[205] P. Huber, A. M. Jensen, L. O. Jepsen, and K. Jensen. Reachability trees for high-level Petri nets. Theoretical Computer Science, 45(3):261–292, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ait-Kaci:1986:ASA

[206] H. Ait-Kaci. An algebraic semantics approach to the effective resolution of type equations. Theoretical Computer Science, 45(3):293–351, ???? 1986. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bernot:1986:ADT

[207] G. Bernot, M. Bidoit, and C. Choppy. Abstract data types with exception handling: an initial approach based on a distinction between exceptions and errors. Theoretical Computer Science, 46 (1):13–45, ???? 1986. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Guo:1986:DDL

[208] Y. Q. Guo, G. W. Xu, and G. Thierrin. Disjunctive decomposition of languages. Theoretical Computer Science, 46(1):47–51, ???? 1986. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hashiguchi:1986:NFG

[209] K. Hashiguchi. Notes on finitely generated semigroups and pumping conditions for regular languages. Theoretical Computer Science, 46(1):53–66, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lerat:1986:NN

[210] N. Lerat and W. Lipski, Jr. Nonapplicable nulls. Theoretical Computer Science, 46(1):67–82, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Head:1986:PDL

[211] T. Head and B. Lando. Periodic D0L languages. Theoretical Computer Science, 46(1):83–89, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yamasaki:1986:COR

[212] H. Yamasaki, M. Takahashi, and K. Kobayashi. Characterization of omegaregular languages by monadic secondorder formulas. Theoretical Computer Science, 46(1):91–99, ???? 1986. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Latteux:1986:TCR

[213] M. Latteux and E. Timmerman. Two characterizations of rational adherences. Theoretical Computer Science, 46(1): 101–106, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Howell:1986:SCB

[214] R. R. Howell, L. E. Rosier, Dung T.

Huynh, and Hsu-Chun Yen. Some complexity bounds for problems concerning finite and 2- dimensional vector addition systems with states. Theoretical Computer Science, 46(2–3):107– 140, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jaffar:1986:SIT

[215] J. Jaffar and P. J. Stuckey. Semantics of infinite tree logic programming. Theoretical Computer Science, 46(2–3): 141–158, ???? 1986. CODEN TCSCDI.

ISSN 0304-3975 (print), 1879-2294 (electronic).

Duboc:1986:SEF

[216] C. Duboc. On some equations in free partially commutative monoids. Theoretical Computer Science, 46(2–3):159– 174, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brazma:1986:GRE

[217] A. Brazma and E. Kinber. Generalized regular expressions — a language for synthesis of programs with branching in loops. Theoretical Computer Science, 46(2–3):175–195, ???? 1986. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Longo:1986:CHT

[218] G. Longo and S. Martini. Computability in higher types, P omega and the completeness of type assignment. Theoretical Computer Science, 46(2–3):197– 217, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dieu:1986:APT

[219] Phan Dinh Dieu, Le Cong Thanh, and Le Tuan Hoa. Average polynomial time complexity of some NP-complete problems. Theoretical Computer Science, 46 (2–3):219–237, ???? 1986. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hajek:1986:SDL

[220] P. Hajek. A simple dynamic logic. Theoretical Computer Science, 46(2–3):239– 259, ???? 1986. CODEN TCSCDI.

ISSN 0304-3975 (print), 1879-2294 (electronic).

Huang:1986:APT

[221] Chua-Huang Huang and C. Lengauer.

The automated proof of a trace transformation for a bitonic sort. Theoretical Computer Science, 46(2–3):261– 284, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dicky:1986:AAM

[222] A. Dicky. An algebraic and algorithmic method for analysing transition systems. Theoretical Computer Science, 46(2–3): 285–303, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hardin:1986:PTR

[223] T. Hardin and A. Laville. Proof of termination of the rewriting system SUBST on CCL. Theoretical Computer Science, 46(2–3):305–312, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diekert:1986:SVE

[224] V. Diekert. On some variants of the Ehrenfeucht Conjecture. Theoretical Computer Science, 46(2–3):313– 318, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diekert:1986:CMC

[225] V. Diekert. Commutative monoids have complete presentations by free (non-commutative) monoids. Theoretical Computer Science, 46(2–3):319– 327, ???? 1986. CODEN TCSCDI.

ISSN 0304-3975 (print), 1879-2294 (electronic).

Griesser:1986:LBA

[226] B. Griesser. Lower bounds for the approximative complexity. Theoretical Computer Science, 46(2–3):329– 338, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Esik:1986:CCA

[227] Z. Esik and P. Domosi. Complete classes of automata for the α_{0}-product. Theoretical Computer Science, 47(1):1– 14, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Culik:1986:RTP

[228] K. Culik, II and Sheng Yu. Real-time, pseudo real-time, and linear-time ITA. Theoretical Computer Science, 47(1):15– 26, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Narendran:1986:PCE

[229] P. Narendran and F. Otto. The problems of cyclic equality and conjugacy for finite complete rewriting systems. Theoretical Computer Science, 47(1):27–38, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Johnson:1986:RER

[230] J. H. Johnson. Rational equivalence relations. Theoretical Computer Science, 47(1):39–60, ???? 1986. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pelc:1986:LPS

[231] A. Pelc. Lie patterns in search procedures. Theoretical Computer Science, 47(1):61–69, ???? 1986. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Culik:1986:EFV

[232] K. Culik, II and J. Karhumaki. The equivalence of finite valued transducers (on HDT0L languages) is decidable. Theoretical Computer Science, 47(1):71– 84, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Valiant:1986:NED

[233] L. G. Valiant and V. V. Vazirani. NP is as easy as detecting unique solutions. Theoretical Computer Science, 47(1):85– 93, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pecuchet:1986:CBA

[234] J.-P. Pecuchet. On the complementation of Buchi Automata. Theoretical Computer Science, 47(1):95–98, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Heintz:1986:PSG

[235] J. Heintz. On polynomials with symmetric Galois group which are easy to compute. Theoretical Computer Science, 47 (1):99–105, ???? 1986. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wyrostek:1986:SUC

[236] P. Wyrostek. On the size of unambiguous context-free grammars. Theoretical Computer Science, 47(1):107–110, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dymond:1986:NPC

[237] P. W. Dymond. On nondeterminism in parallel computation. Theoretical Computer Science, 47(2):111–120, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Orponen:1986:CCC

[238] P. Orponen. A classification of complexity core lattices. Theoretical Computer Science, 47(2):121–130, ???? 1986. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wagner:1986:SOC

[239] K. W. Wagner. Some observations on the connection between counting and recursion. Theoretical Computer Science, 47(2):131–147, ???? 1986. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chrobak:1986:FAU

[240] M. Chrobak. Finite automata and unary languages. Theoretical Computer Science, 47(2):149–158, ???? 1986. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See errata [634].

Dare:1986:ST

[241] V. R. Dare and R. Siromoney. Subword topology. Theoretical Computer Science, 47(2):159–168, ???? 1986. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Homer:1986:SCS

[242] S. Homer. On simple and creative sets in NP. Theoretical Computer Science, 47 (2):169–180, ???? 1986. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Spehner:1986:RCS

[243] J.-C. Spehner. The rapid calculation of shuffles of two words. Theoretical Computer Science, 47(2):191–203, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kirousis:1986:SP

[244] L. M. Kirousis and C. H. Papadimitriou.

Searching and pebbling. Theoretical Computer Science, 47(2):205–218, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Parchmann:1986:SEI

[245] R. Parchmann and J. Duske. Selfembedding indexed grammars. Theoretical Computer Science, 47(2):219– 223, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Otto:1986:USE

[246] F. Otto. The undecidability of selfembedding for finite semi-Thue and Thue systems. Theoretical Computer Science, 47(2):225–232, ???? 1986. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Karchmer:1986:TTS

[247] M. Karchmer. Two time-space tradeoffs for element distinctness. Theoretical Computer Science, 47(3):237–246, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Maon:1986:BMV

[248] Y. Maon and A. Yehudai. Balance of many-valued transductions and equivalence problems. Theoretical Computer Science, 47(3):247–262, ???? 1986. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ko:1986:OWF

[249] Ker-I Ko, T. J. Long, and D.-Z. Du. On one-way functions and polynomial-time isomorphisms. Theoretical Computer Science, 47(3):263–276, ???? 1986. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Maon:1986:PED

[250] Y. Maon, B. Schieber, and U. Vishkin. Parallel ear decomposition search (EDS) and st-numbering in graphs. Theoretical Computer Science, 47(3):277–298, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ko:1986:CFR

[251] Ker-I Ko. On the continued fraction representation of computable real numbers. Theoretical Computer Science, 47 (3):299–313, ???? 1986. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rytter:1986:CPP

[252] W. Rytter. On the complexity of parallel parsing of general context-free languages. Theoretical Computer Science, 47(3):315–321, ???? 1986. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hong:1986:LBC

[253] J. W. Hong and Q. Zuo. Lower bounds on communication overlap of networks. Theoretical Computer Science, 47(3): 323–328, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Szalas:1986:CSC

[254] Z. Szalas. Concerning the semantic consequence relation in first-order temporal logic. Theoretical Computer Science, 47 (3):329–334, ???? 1986. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Deleage:1986:RID

[255] J.-L. Deleage and L. Pierre. The rational index of the Dyck language D′_{1}∗. Theoretical Computer Science, 47(3):335– 343, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Esik:1986:PP

[256] Z. Esik and F. Gecseg. On α_{0}-products and α_{2}-products. Theoretical Computer Science, 48(1):1–8, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ko:1986:NIP

[257] K.-I. Ko. On the notion of infinite pseudorandom sequences. Theoretical Computer Science, 48(1):9–33, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Spehner:1986:OFG

[258] J.-C. Spehner. The occurrence of the factors of a given word in a text.

Theoretical Computer Science, 48(1):35– 52, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bublitz:1986:PCM

[259] S. Bublitz, U. Schurfeld, I. Wegener, and B. Voigt. Properties of complexity measures for PRAMs and WRAMs. Theoretical Computer Science, 48(1):53– 73, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kruskal:1986:UTI

[260] C. P. Kruskal and M. Snir. A unified theory of interconnection network structure. Theoretical Computer Science, 48 (1):75–94, ???? 1986. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Statman:1986:ECP

[261] R. Statman. Every countable poset is embeddable in the poset of unsolvable terms. Theoretical Computer Science, 48 (1):95–100, ???? 1986. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Head:1986:RSI

[262] T. Head and B. Lando. Regularity of sets of initial strings of periodic D0Lsystems. Theoretical Computer Science, 48(1):101–108, ???? 1986. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hromkovic:1986:CCH

[263] J. Hromkovic. Communication complexity hierarchy. Theoretical Computer Science, 48(1):109–115, ???? 1986. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berry:1986:RED

[264] G. Berry and R. Sethi. From regular expressions to deterministic automata. Theoretical Computer Science, 48(1): 117–126, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Weispfenning:1986:CWP

[265] V. Weispfenning. The complexity of the word problem for Abelian l-groups. Theoretical Computer Science, 48(1):127– 132, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tchuente:1986:SSP

[266] M. Tchuente. Sequential simulation of parallel iterations and applications. Theoretical Computer Science, 48(2–3):135– 144, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Marek:1986:ASE

[267] W. Marek and H. Rasiowa. Approximating sets with equivalence relations. Theoretical Computer Science, 48(2–3): 145–152, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chrobak:1986:HOW

[268] M. Chrobak. Hierarchies of one-way multihead automata languages. Theoretical Computer Science, 48(2–3):153– 181, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Duboc:1986:MPA

[269] C. Duboc. Mixed product and asynchronous automata. Theoretical Computer Science, 48(2–3):183–199, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ehrenfeucht:1986:AFU

[270] A. Ehrenfeucht, H. J. Hoogeboom, and G. Rozenberg. On the active and full use of memory in right-boundary grammars and push-down automata. Theoretical Computer Science, 48(2–3):201– 228, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fitting:1986:PML

[271] M. Fitting. Partial models and logic programming. Theoretical Computer Science, 48(2–3):229–255, ???? 1986. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Danko:1986:FOA

[272] W. Danko. First-order approximation of algorithmic theories. Theoretical Computer Science, 48(2–3):257–272, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Italiano:1986:AEP

[273] G. F. Italiano. Amortized efficiency of a path retrieval data structure. Theoretical Computer Science, 48(2–3):273– 281, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Salomaa:1986:PKC

[274] A. Salomaa and S. Yu. On a publickey cryptosystem based on iterated morphisms and substitutions. Theoretical Computer Science, 48(2–3):283– 296, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ginsburg:1986:POH

[275] S. Ginsburg and C. Tang. Projection of object histories. Theoretical Computer Science, 48(2–3):297–328, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gibbons:1986:DSP

[276] A. Gibbons and W. Rytter. On the decidability of some problems about rational subsets of free partially commutative monoids. Theoretical Computer Science, 48(2–3):329–337, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Takahashi:1987:BHO

[277] M. Takahashi. Brzozowski hierarchy of omega-languages. Theoretical Computer Science, 49(1):1–12, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Squier:1987:USC

[278] C. C. Squier. Units of special Church– Rosser monoids. Theoretical Computer Science, 49(1):13–22, ???? 1987. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Parchmann:1987:GDM

[279] R. Parchmann and J. Duske. Grammars, derivation modes and properties of indexed and type-0 languages. Theoretical Computer Science, 49(1):23–42, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Oyamaguchi:1987:CRP

[280] M. Oyamaguchi. The Church–Rosser property for ground term-rewriting systems is decidable. Theoretical Computer Science, 49(1):43–79, ???? 1987. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Guibas:1987:SRT

[281] L. J. Guibas, J. Stolfi, and K. L. Clarkson. Solving related two- and threedimensional linear programming problems in logarithmic time. Theoretical Computer Science, 49(1):81–84, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:1987:TIC

[282] Anonymous. Twelfth International Colloquium on Automata, Languages and Programming. Theoretical Computer Science, 49(2–3):??, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deBakker:1987:ISF

[283] J. W. de Bakker, J.-J. C. Meyer, and E.-R. Olderog. Infinite streams and finite observations in the semantics of uniform concurrency. Theoretical Computer Science, 49(2–3):87–112, 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Main:1987:AIS

[284] M. G. Main, W. Bucher, and D. Haussler. Applications of an infinite squarefree co-CFL. Theoretical Computer Science, 49(2–3):113–119, 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hennessy:1987:ATF

[285] M. Hennessy. An algebraic theory of fair asynchronous communicating processes. Theoretical Computer Science, 49 (2–3):121–143, 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bouge:1987:RSD

[286] L. Bouge. Repeated snapshots in distributed systems with synchronous communications and their implementation in CSP. Theoretical Computer Science, 49(2–3):145–169, 1987. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Galil:1987:DAS

[287] Z. Galil, G. M. Landau, and M. M.

Yung. Distributed algorithms in synchronous broadcasting networks. Theoretical Computer Science, 49(2–3): 171–184, 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Larsen:1987:CDE

[288] K. G. Larsen. A context dependent equivalence between processes. Theoretical Computer Science, 49(2–3):185–215, 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sistla:1987:CPB

[289] A. Prasad Sistla, M. Y. Vardi, and P. Wolper. The complementation problem for Buchi automata with applications to temporal logic. Theoretical Computer Science, 49(2–3):217–237, 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cole:1987:PPS

[290] R. Cole. Partitioning point sets in arbitrary dimension. Theoretical Computer Science, 49(2–3):239–265, 1987. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kurtz:1987:PDA

[291] T. G. Kurtz and U. Manber. A probabilistic distributed algorithm for set intersection and its analysis. Theoretical Computer Science, 49(2–3):267–282, 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Flajolet:1987:AMA

[292] P. Flajolet. Analytic models and ambiguity of context-free languages. Theoretical Computer Science, 49(2–3):283–309, 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stirling:1987:MLC

[293] C. Stirling. Modal logics for communicating systems. Theoretical Computer Science, 49(2–3):311–347, 1987. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Girard:1987:LL

[294] J.-Y. Girard. Linear logic. Theoretical Computer Science, 50(1):1–102, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gray:1987:CAD

[295] J. W. Gray. Categorical aspects of data type constructors. Theoretical Computer Science, 50(2):103–135, ???? 1987. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bergstra:1987:ASC

[296] J. A. Bergstra and J. V. Tucker. Algebraic specifications of computable and semicomputable data types. Theoretical Computer Science, 50(2):137–181, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mazoyer:1987:SSM

[297] J. Mazoyer. A six-state minimal time solution to the firing squad synchronization problem. Theoretical Computer Science, 50(2):183–238, ???? 1987. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Phillips:1987:RT

[298] I. Phillips. Refusal testing. Theoretical Computer Science, 50(3):241–284, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sain:1987:TCN

[299] I. Sain. Total correctness in nonstandard logics of programs. Theoretical Computer Science, 50(3):285–321, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wagner:1987:CCP

[300] E. G. Wagner and H. Ehrig. Canonical constraints for parameterized data types. Theoretical Computer Science, 50 (3):323–349, ???? 1987. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Troelstra:1987:SML

[301] A. S. Troelstra. On the syntax of Martin-Löf’s type theories. Theoretical Computer Science, 51(1–2):1–26, ????

1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

LeChenadec:1987:ADA

[302] P. Le Chenadec. Analysis of Dehn’s algorithm by critical pairs. Theoretical Computer Science, 51(1–2):27–52, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wagner:1987:MCQ

[303] K. W. Wagner. More complicated questions about maxima and minima, and some closures of NP. Theoretical Computer Science, 51(1–2):53–80, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Habel:1987:CGL

[304] A. Habel and H. J. Kreowski. Characteristics of graph languages generated by edge replacement. Theoretical Computer Science, 51(1–2):81–115, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beauquier:1987:NCT

[305] J. Beauquier and F. Gire. A note on the characterisation theorem for algebraic generators. Theoretical Computer Science, 51(1–2):117–127, ???? 1987. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baeten:1987:CKF

[306] J. C. M. Baeten, J. A. Bergstra, and J. W. Klop. On the consistency of Koomen’s fair abstraction rule. Theoretical Computer Science, 51(1–2):129– 176, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ambos-Spies:1987:DPT

[307] K. Ambos-Spies, H. Fleischhack, and H. Huwig. Diagonalizations over polynomial time computable sets. Theoretical Computer Science, 51(1–2):177– 204, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Orejas:1987:CPC

[308] F. Orejas. A characterization of passing compatibility for parameterized specifications. Theoretical Computer Science, 51(1–2):205–214, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Carpi:1987:URM

[309] A. Carpi. On unambiguous reductions of monoids of unambiguous relations. Theoretical Computer Science, 51(1–2):215– 220, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vyskoc:1987:TSA

[310] J. Vyskoc. An O(n^{lgk}.2^{n}2) time and O(k.2^{n}k) space algorithm for certain NP-complete problems. Theoretical Computer Science, 51(1–2):221– 227, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Belaga:1987:CUA

[311] E. G. Belaga. Constructive universal algebra: an introduction. Theoretical Computer Science, 51(1–2):229– 238, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Parberry:1987:TRS

[312] I. Parberry. On the time required to sum n semigroup elements on a parallel machine with simultaneous writes. Theoretical Computer Science, 51(1–2):239– 247, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Head:1987:BDL

[313] T. Head and B. Lando. Bounded DL languages. Theoretical Computer Science, 51(3):255–264, ???? 1987. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Homer:1987:HPD

[314] S. Homer and T. J. Long. Honest polynomial degrees and P=?NP. Theoretical Computer Science, 51(3):265–280, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berard:1987:LSF

[315] B. Bérard. Literal shuffle (formal languages). Theoretical Computer Science, 51(3):281–299, ???? 1987. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yokomori:1987:PMC

[316] T. Yokomori. On purely morphic characterizations of context-free languages. Theoretical Computer Science, 51(3): 301–308, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yamasaki:1987:FSH

[317] S. Yamasaki, M. Yoshida, and S. Doshita. A fixpoint semantics of Horn sentences based on substitution sets. Theoretical Computer Science, 51(3):309–324, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hromkovic:1987:RBN

[318] J. Hromkovic. Reversal-bounded nondeterministic multicounter machines and complementation. Theoretical Computer Science, 51(3):325–330, ???? 1987. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beth:1987:CCG

[319] T. Beth. On the computational complexity of the general discrete Fourier transform. Theoretical Computer Science, 51 (3):331–339, ???? 1987. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Galil:1987:PSM

[320] Z. Galil and R. Giancarlo. Parallel string matching with k mismatches. Theoretical Computer Science, 51(3):341– 348, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zaionc:1987:WOD

[321] M. Zaionc. Word operation definable in the typed lambda-calculus. Theoretical Computer Science, 52(1–2):1–14, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ko:1987:HRO

[322] K. I. Ko. On helping by robust oracle machines. Theoretical Computer Science, 52(1–2):15–36, ???? 1987. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kennaway:1987:GR

[323] R. Kennaway. On ’on graph rewritings’. Theoretical Computer Science, 52(1–2): 37–58, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sakarovitch:1987:RTL

[324] J. Sakarovitch. On regular trace languages. Theoretical Computer Science, 52(1–2):59–75, ???? 1987. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vonzurGathen:1987:FPP

[325] J. von zur Gathen. Factoring polynomials and primitive elements for special primes. Theoretical Computer Science, 52(1–2):77–89, ???? 1987. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Seite:1987:YEL

[326] B. Seite. A Yacc extension for LRR grammar parsing. Theoretical Computer Science, 52(1–2):91–143, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mundici:1987:SMV

[327] D. Mundici. Satisfiability in manyvalued sentential logic is NP-complete. Theoretical Computer Science, 52(1–2): 145–153, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Spirakis:1987:PCD

[328] P. Spirakis. The parallel complexity of deadlock detection. Theoretical Computer Science, 52(1–2):155–163, ????

1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Moriya:1987:TCI

[329] T. Moriya. Topological characterizations of infinite tree languages. Theoretical Computer Science, 52(1–2):165– 171, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kim:1987:MEP

[330] C. Kim and I. H. Sudborough. The membership and equivalence problems for picture languages. Theoretical Computer Science, 52(3):177–191, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Marberg:1987:DSA

[331] J. M. Marberg and E. Gafni. Distributed sorting algorithms for multichannel broadcast networks. Theoretical Computer Science, 52(3):193–203, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Felleisen:1987:STS

[332] M. Felleisen, D. P. Friedman, E. Kohlbecker, and B. Duba. A syntactic theory of sequential control. Theoretical Computer Science, 52(3):205–237, ???? 1987. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nigmatullin:1987:MLB

[333] R. G. Nigmatullin. Models of lowerbounds proofs. Theoretical Computer Science, 52(3):239–249, ???? 1987. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Balcazar:1987:CCP

[334] J. L. Balcazar, J. Diaz, and J. Gabarro. On characterizations of the class PSPACE/ poly. Theoretical Computer Science, 52(3):251–267, ???? 1987. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gelenbe:1987:SDF

[335] E. Gelenbe and D. Finkel. Stationary deterministic flows. II. the single-server queue. Theoretical Computer Science, 52 (3):269–280, ???? 1987. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Topor:1987:DIF

[336] R. W. Topor. Domain-independent formulas and databases. Theoretical Computer Science, 52(3):281–306, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stefanescu:1987:FTI

[337] G. Stefanescu. On flowchart theories. II. the nondeterministic case. Theoretical Computer Science, 52(3):307–340, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brandenburg:1987:CDA

[338] F. J. Brandenburg. Comments on ’deque automata and a subfamily of contextsensitive languages which contains all semilinear bounded languages’. Theoretical Computer Science, 52(3):341– 342, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:1987:ECT

[339] Anonymous. Eleventh colloquium on trees in algebra and programming. Theoretical Computer Science, 53(1): ??, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wagner:1987:CTP

[340] E. G. Wagner. A categorical treatment of pre- and post-conditions. Theoretical Computer Science, 53(1):3–24, 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

File:1987:CIA

[341] G. File. Classical and incremental attribute evaluation by means of recursive procedures. Theoretical Computer Science, 53(1):25–65, 1987. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

FrutosEscrig:1987:PIS

[342] D. Frutos Escrig. Probabilistic Ianov’s schemes. Theoretical Computer Science, 53(1):67–97, 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Louchard:1987:RWG

[343] G. Louchard. Random walks, Gaussian processes and list structures. Theoretical Computer Science, 53(1):99–124, 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pelin:1987:BEC

[344] A. Pelin and J. H. Gallier. Building exact computation sequences. Theoretical Computer Science, 53(1):125–150, 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pettorossi:1987:DEP

[345] A. Pettorossi. Derivation of efficient programs for computing sequences of actions. Theoretical Computer Science, 53 (1):151–167, 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Attiya:1987:LCS

[346] H. Attiya and Y. Mansour. Language complexity on the synchronous anonymous ring. Theoretical Computer Science, 53(2–3):169–185, ???? 1987. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Litovsky:1987:GRO

[347] I. Litovsky and E. Timmerman. On generators of rational omega-power languages. Theoretical Computer Science, 53(2–3):187–200, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schnorr:1987:HPT

[348] C. P. Schnorr. A hierarchy of polynomial time lattice basis reduction algorithms. Theoretical Computer Science, 53(2–3): 201–224, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abramsky:1987:OET

[349] S. Abramsky. Observation equivalence as a testing equivalence. Theoretical Computer Science, 53(2–3):225– 241, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Esik:1987:RTA

[350] Z. Esik and F. Gecseg. On a representation of tree automata. Theoretical Computer Science, 53(2–3):243– 255, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Muller:1987:NCS

[351] H. Muller and A. Brandstadt. The NPcompleteness of Steiner tree and dominating set for chordal bipartite graphs. Theoretical Computer Science, 53(2–3): 257–265, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beynon:1987:PMC

[352] M. Beynon and J. Buckle. On the planar monotone computation of Boolean functions. Theoretical Computer Science, 53 (2–3):267–279, ???? 1987. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Peleg:1987:GPR

[353] D. Peleg and E. Upfal. The generalized packet routing problem. Theoretical Computer Science, 53(2–3):281– 293, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rytter:1987:OPP

[354] W. Rytter and R. Giancarlo. Optimal parallel parsing of bracket languages. Theoretical Computer Science, 53(2–3): 295–306, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Garzon:1987:CA

[355] M. Garzon. Cyclic automata. Theoretical Computer Science, 53(2–3):307– 317, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nishida:1987:RWS

[356] T. Nishida and Y. Kobuchi. Repeatable words for substitution. Theoretical Computer Science, 53(2–3):319– 333, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hallnas:1987:ICL

[357] L. Hallnas. An intensional characterization of the largest bisimulation. Theoretical Computer Science, 53(2–3):335– 343, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Santoro:1987:GCV

[358] N. Santoro, J. B. Sidney, S. J. Sidney, and J. Urrutia. Geometric containment and vector dominance. Theoretical Computer Science, 53(2–3):345– 352, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hsiang:1987:AIT

[359] Jieh Hsiang and M. Srivas. Automatic inductive theorem proving using PRO-LOG. Theoretical Computer Science, 54 (1):3–28, September 1987. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jensen:1987:RDD

[360] F. V. Jensen and K. D. Larsen. Recursively defined domains and their induction principles. Theoretical Computer Science, 54(1):29–51, September 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Marchetti-Spaccamela:1987:NPE

[361] A. Marchetti-Spaccamela. New protocols for the election of a leader in a ring. Theoretical Computer Science, 54(1):53– 64, September 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pan:1987:CPM

[362] V. Pan. Complexity of parallel matrix computations. Theoretical Computer Science, 54(1):65–85, September 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bajaj:1987:GOP

[363] C. Bajaj. Geometric optimization and the polynomial hierarchy. Theoretical Computer Science, 54(1):87–102, September 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lakshmanan:1987:ATF

[364] V. S. Lakshmanan and C. E. Veni Madhavan. An algebraic theory of functional and multivalued dependencies in relational databases. Theoretical Computer Science, 54(1):103–128, September 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Venkatesan:1987:AAW

[365] S. M. Venkatesan. Approximation algorithms for weighted matching. Theoretical Computer Science, 54(1):129– 137, September 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Priese:1987:IRT

[366] L. Priese, R. Rehrmann, and U. Willecke-Klemme. An introduction to the regular theory of fairness. Theoretical Computer Science, 54(2–3):139–163, October 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rindone:1987:CFC

[367] G. Rindone. Construction of a family of codes associated with certain finite groups. Theoretical Computer Science, 54(2–3):165–179, October 1987. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brandstadt:1987:DPP

[368] A. Brandstadt and D. Kratsch. On domination problems for permutation and other graphs. Theoretical Computer Science, 54(2–3):181–198, October 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Szalas:1987:CAC

[369] A. Szalas. A complete axiomatic characterization of first-order temporal logic of linear time. Theoretical Computer Science, 54(2–3):199–214, October 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cockett:1987:DDT

[370] J. R. B. Cockett. Discrete decision theory: manipulations. Theoretical Computer Science, 54(2–3):215–236, October 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bertolazzi:1987:MAR

[371] P. Bertolazzi and A. Sassano. An O(mn)

algorithm for regular set-covering problems. Theoretical Computer Science, 54 (2–3):237–247, October 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Watanabe:1987:CPT

[372] O. Watanabe. A comparison of polynomial time completeness notions. Theoretical Computer Science, 54(2–3):249– 265, October 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Muller:1987:AAI

[373] D. E. Muller and P. E. Schupp. Alternating automata on infinite trees. Theoretical Computer Science, 54(2–3):267– 276, October 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fauconnier:1987:ASI

[374] H. Fauconnier. Asynchronous semantics and infinite behaviour in CSP. Theoretical Computer Science, 54(2–3):277– 298, October 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ginsburg:1987:CFI

[375] S. Ginsburg and Chang jie Tang. Canonical forms for interval functions. Theoretical Computer Science, 54(2–3):299– 313, October 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sado:1987:FEC

[376] K. Sado and Y. Igarashi. A function for evaluating the computing time of a bubbling system. Theoretical Computer Science, 54(2–3):315–324, October 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Iwata:1987:SPT

[377] S. Iwata and T. Kasai. Simultaneous (poly-time, log-space) lower bounds. Theoretical Computer Science, 54(2–3): 325–329, October 1987. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liskiewicz:1987:RBA

[378] M. Liskiewicz, K. Lorys, and M. Piotrow. On reversal bounded alternating Turing machines. Theoretical Computer Science, 54(2–3):331–339, October 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ko:1987:CFR

[379] K.-I. Ko. On the continued fraction representation of computable real numbers. Theoretical Computer Science, 54(2–3): 341–343, October 1987. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Peleg:1987:CPS

[380] D. Peleg. Concurrent program schemes and their logics. Theoretical Computer Science, 55(1):1–45, November 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Seidl:1987:PRH

[381] H. Seidl. Parameter-reduction of higher level grammars. Theoretical Computer Science, 55(1):47–85, November 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Best:1987:SCB

[382] E. Best and R. Devillers. Sequential and concurrent behaviour in Petri net theory. Theoretical Computer Science, 55 (1):87–136, November 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Courcelle:1987:ADC

[383] B. Courcelle. An axiomatic definition of context-free rewriting and its application to NLC graph grammars. Theoretical Computer Science, 55(2–3):141– 181, December 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brandenburg:1987:RLF

[384] F.-J. Brandenburg. Representations of language families by homomorphic equality operations and generalized equality sets. Theoretical Computer Science, 55(2–3):183–263, December 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bartha:1987:EAS

[385] M. Bartha. An equational axiomatization of systolic systems. Theoretical Computer Science, 55(2–3):265–289, December 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lehmann:1987:UMC

[386] J.-P. Lehmann. A universal machine without change of state. Theoretical Computer Science, 55(2–3):291–348, December 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Armbruster:1988:PDM

[387] D. Armbruster. A polynomial determination of the most-recent property in Pascal-like programs. Theoretical Computer Science, 56(1):3–15, January 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hankin:1988:SAP

[388] C. Hankin, G. Burn, and S. Peyton Jones. A safe approach to parallel combinator reduction. Theoretical Computer Science, 56(1):17–36, January 1988. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kaplan:1988:RNC

[389] S. Kaplan. Rewriting with a nondeterministic choice operator. Theoretical Computer Science, 56(1):37–57, January 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nielson:1988:TLS

[390] F. Nielson and H. R. Nielson. Twolevel semantics and code generation. Theoretical Computer Science, 56(1):59– 133, January 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stark:1988:PEB

[391] E. W. Stark. Proving entailment between conceptual state specifications.

Theoretical Computer Science, 56(1): 135–154, January 1988. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fachini:1988:CTS

[392] E. Fachini and M. Napoli. C-tree systolic automata. Theoretical Computer Science, 56(2):155–186, February 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Marek:1988:NSM

[393] W. Marek. A natural semantics for modal logic over databases. Theoretical Computer Science, 56(2):187–209, February 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Iliopoulos:1988:CCA

[394] C. S. Iliopoulos. On the computational complexity of the Abelian permutation group structure, membership and intersection problems. Theoretical Computer Science, 56(2):211–222, February 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sekimoto:1988:OSR

[395] S. Sekimoto and S. Hirokawa. One-step recurrent terms in λ-β-calculus. Theoretical Computer Science, 56(2):223– 231, February 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Carpi:1988:MUC

[396] A. Carpi. Multidimensional unrepetitive configurations. Theoretical Computer Science, 56(2):233–241, February 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mohrherr:1988:RLP

[397] J. Mohrherr. A remark on the length problem. Theoretical Computer Science, 56(2):243–248, February 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tretkoff:1988:CCG

[398] C. Tretkoff. Complexity, combinatorial group theory and the language of palutators. Theoretical Computer Science, 56 (3):253–275, March 1988. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Staples:1988:DUA

[399] J. Staples. Delaying unification algorithms for lambda calculi. Theoretical Computer Science, 56(3):277– 288, March 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gradel:1988:SPA

[400] E. Gradel. Subclasses of Presburger arithmetic and the polynomial-time hierarchy. Theoretical Computer Science, 56(3):289–301, March 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Martino:1988:QIG

[401] A. Martino. A quantitative interpretation of Girard’s system F. Theoretical Computer Science, 56(3):303– 320, March 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Boonyavatana:1988:IPL

[402] R. Boonyavatana and G. Slutzki. The interchange or pump (DI)lemmas for context-free languages. Theoretical Computer Science, 56(3):321– 338, March 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bruyere:1988:AQA

[403] V. Bruyere. An answer to a question about finite maximal prefix sets of words. Theoretical Computer Science, 56 (3):339–344, March 1988. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baader:1988:UCI

[404] F. Baader and W. Buttner. Unification in commutative idempotent monoids. Theoretical Computer Science, 56(3): 345–353, March 1988. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Broy:1988:ESP

[405] M. Broy. Equational specification of partial higher-order algebras. Theoretical Computer Science, 57(1):3–45, April 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ibarra:1988:TDI

[406] O. H. Ibarra and M. A. Palis. Twodimensional iterative arrays: characterizations and applications. Theoretical Computer Science, 57(1):47–86, April 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ablyv:1988:CPP

[407] F. M. Ablyv. The complexity properties of probabilistic automata with isolated cut point. Theoretical Computer Science, 57(1):87–95, April 1988. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hromkovic:1988:ANA

[408] J. Hromkovic. The advantages of a new approach to defining the communication complexity for VLSI. Theoretical Computer Science, 57(1):97–111, April 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jukna:1988:ECC

[409] S. P. Jukna. Entropy of contact circuits and lower bounds on their complexity. Theoretical Computer Science, 57(1):113–129, April 1988. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tarhio:1988:GAA

[410] Jorma Tarhio and Esko Ukkonen. A greedy approximation algorithm for constructing shortest common superstrings. Theoretical Computer Science, 57(1): 131–145, April 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kuich:1988:MSP

[411] W. Kuich. Matrix systems and principal cones of algebraic power series. Theoretical Computer Science, 57(1):147– 152, April 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Skarbek:1988:GOT

[412] Wladyslaw Skarbek. Generating ordered trees. Theoretical Computer Science, 57 (1):153–159, April 1988. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Avron:1988:SPT

[413] A. Avron. The semantics and proof theory of linear logic. Theoretical Computer Science, 57(2–3):161–184, May 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pierre:1988:RIC

[414] L. Pierre and J.-M. Farinone. Rational index of context-free languages in exp(Θ( ^{p} \sqrt{n})) and n^{Θ((ln n)^{(}1/p))}. Theoretical Computer Science, 57(2–3):185– 204, May 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Velinov:1988:ASD

[415] Y. Velinov. An algebraic structure for derivations in rewriting systems. Theoretical Computer Science, 57(2–3):205– 224, May 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ibarra:1988:RPC

[416] O. H. Ibarra and Tao Jiang. Relating the power of cellular arrays to their closure properties. Theoretical Computer Science, 57(2–3):225–238, May 1988. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Duchamp:1988:TTP

[417] G. Duchamp and J. Y. Thibon. Transfer theorems for partially commutative polynomials. Theoretical Computer Science, 57(2–3):239–249, May 1988. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Meyer:1988:ACS

[418] J.-J. C. Meyer and E. P. de Vink. Applications of compactness in the Smyth powerdomain of streams. Theoretical Computer Science, 57(2–3):251– 282, May 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beal:1988:CCF

[419] M. P. Beal. Circular codes, finite automata and entropy. Theoretical Computer Science, 57(2–3):283–302, May 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hashiguchi:1988:NCR

[420] K. Hashiguchi. Notes on congruence relations and factor pumping conditions for rational languages. Theoretical Computer Science, 57(2–3):303–316, May 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Szalas:1988:IFO

[421] A. Szalas and L. Holenderski. Incompleteness of first-order temporal logic with until. Theoretical Computer Science, 57(2–3):317–325, May 1988. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Grossman:1988:IIC

[422] J. W. Grossman and R. S. Zeitman. An inherently iterative computation of Ackermann’s function. Theoretical Computer Science, 57(2–3):327–330, May 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Arques:1988:CAC

[423] D. Arques, J. Francon, M. T. Guichet, and P. Guichet. Comparison of algorithms controlling concurrent access to a database: A combinatorial approach. Theoretical Computer Science, 58(1-3): 3–16, June 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Averbuch:1988:CAM

[424] Amir Averbuch, Zvi Galil, and Shmuel Winograd. Classification of all the minimal bilinear algorithms for computing the coefficients of the product of two polynomials modulo a polynomial. Part I. the algebra G[u]/ < Q(u)^{ℓ} >, ℓ > 1. Theoretical Computer Science, 58(1-3): 17–56, June 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Borodin:1988:TBS

[425] Allan Borodin, Faith E. Fich, Friedhelm Meyer auf der Heide, Eli Upfal, and Avi Wigderson. A tradeoff between search and update time for the implicit dictionary problem. Theoretical Computer Science, 58(1-3):57–68, June 1988. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brandenburg:1988:ISQ

[426] Franz J. Brandenburg. On the intersection of stacks and queues. Theoretical Computer Science, 58(1-3):69–80, June 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Choffrut:1988:CRF

[427] C. Choffrut and M. P. Schütz en berger.

Counting with rational functions. Theoretical Computer Science, 58(1-3):81– 101, June 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DeFelice:1988:FBS

[428] Clelia De Felice. Finite biprefix sets of paths in a graph. Theoretical Computer Science, 58(1-3):103–128, June 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hartmanis:1988:CCM

[429] Juris Hartmanis and Lane A. Hemachandra. Complexity classes without machines: On complete languages for UP. Theoretical Computer Science, 58 (1-3):129–142, June 1988. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kirschenhofer:1988:FRD

[430] Peter Kirschenhofer and Helmut Prodinger. Further results on digital search trees. Theoretical Computer Science, 58(1-3): 143–154, June 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kraus:1988:KBT

[431] Sarit Kraus and Daniel Lehmann. Knowledge, belief and time. Theoretical Computer Science, 58(1-3):155– 174, June 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lange:1988:DNR

[432] Klaus-Jörn Lange. Decompositions of nondeterministic reductions. Theoretical Computer Science, 58(1-3):175– 181, June 1988. CODEN TCSCDI.

ISSN 0304-3975 (print), 1879-2294 (electronic).

Lisper:1988:SEC

[433] Bjorn Lisper. Synthesis and equivalence of concurrent systems. Theoretical Computer Science, 58(1-3):183– 199, June 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Metivier:1988:RSF

[434] Y. Metivier. On recognizable subsets of free partially commutative monoids. Theoretical Computer Science, 58(1-3): 201–208, June 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Monien:1988:MCN

[435] B. Monien and I. H. Sudborough. Min Cut is NP-complete for edge weighted trees. Theoretical Computer Science, 58 (1-3):209–229, June 1988. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pecuchet:1988:ESP

[436] Jean-Pierre Pecuchet. Etude syntaxique des parties reconnaissables de mots infinis. (French) [Syntactic study of rational omega-languages of infinite words]. Theoretical Computer Science, 58(1-3): 231–248, June 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Reed:1988:TMC

[437] G. M. Reed and A. W. Roscoe. A timed model for communicating sequential processes. Theoretical Computer Science, 58(1-3):249–261, June 1988.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rosier:1988:CDF

[438] Louis E. Rosier and Hsu-Chun Yen. On the complexity of deciding fair termination of probabilistic concurrent finitestate programs. Theoretical Computer Science, 58(1-3):263–324, June 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Simon:1988:IAT

[439] Klaus Simon. An improved algorithm for transitive closure on acyclic digraphs. Theoretical Computer Science, 58(1-3): 325–346, June 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stirling:1988:GOG

[440] Colin Stirling. A generalization of Owicki-Gries’s Hoare logic for a concurrent while language. Theoretical Computer Science, 58(1-3):347–359, June 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Straubing:1988:SLD

[441] Howard Straubing. Semigroups and languages of dot-depth two. Theoretical Computer Science, 58(1-3):361– 378, June 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Varman:1988:EPA

[442] Peter Varman and Kshitij Doshi. An efficient parallel algorithm for updating minimum spanning trees. Theoretical Computer Science, 58(1-3):379– 397, June 1988. CODEN TCSCDI.

ISSN 0304-3975 (print), 1879-2294 (electronic).

Bosco:1988:NVS

[443] Pier Giorgio Bosco, Elio Giovannetti, and Corrado Moiso. Narrowing vs. SLDresolution. Theoretical Computer Science, 59(1–2):3–23, July 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Boudol:1988:CA

[444] G. Boudol and I. Castellani. Concurrency and atomicity. Theoretical Computer Science, 59(1–2):25–84, July 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Breazu-Tannen:1988:EMP

[445] Val Breazu-Tannen and Thierry Coquand. Extensional models for polymorphism. Theoretical Computer Science, 59(1–2):85–114, July 1988. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Browne:1988:CFK

[446] M. C. Browne, E. M. Clarke, and O. Grümberg. Characterizing finite Kripke structures in Propositional Temporal Logic. Theoretical Computer Science, 59(1–2):115–131, July 1988. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Drabent:1988:IAM

[447] Wlodzimierz Drabent and Jan Maluszynski. Inductive assertion method for logic programs. Theoretical Computer Science, 59(1–2):133–155, July 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lafont:1988:LAM

[448] Y. Lafont. The linear abstract machine. Theoretical Computer Science, 59(1–2): 157–180, July 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

RonchiDellaRocca:1988:PTS

[449] Simona Ronchi Della Rocca. Principal type scheme and unification for intersection type discipline. Theoretical Computer Science, 59(1–2):181–209, July 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hesselink:1988:IRU

[450] Wim H. Hesselink. Interpretations of recursion under unbounded nondeterminacy. Theoretical Computer Science, 59 (3):211–234, August 1988. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hesselink:1988:DFM

[451] Wim H. Hesselink. Deadlock and fairness in morphisms of transition systems. Theoretical Computer Science, 59 (3):235–257, August 1988. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hodel:1988:ODE

[452] A. Scottedward Hodel and Michael C. Loui. Optimal dynamic embedding of X-trees into arrays. Theoretical Computer Science, 59(3):259–276, August 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kalorkoti:1988:TIM

[453] K. Kalorkoti. The trace invariant and matrix inversion. Theoretical Computer Science, 59(3):277–286, August 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schmidt-Schauss:1988:ICU

[454] Manfred Schmidt-Schauss. Implication of clauses is undecidable. Theoretical Computer Science, 59(3):287– 296, August 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rytter:1988:EPC

[455] Wojciech Rytter. On efficient parallel computations for some dynamic programming problems. Theoretical Computer Science, 59(3):297–307, August 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Engelfriet:1988:NBN

[456] Joost Engelfriet and George Leih. Nonterminal bounded NLC graph grammars. Theoretical Computer Science, 59 (3):309–315, August 1988. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stoughton:1988:SR

[457] Allen Stoughton. Substitution revisited. Theoretical Computer Science, 59 (3):317–325, August 1988. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aalbersberg:1988:TT

[458] I. J. Aalbersberg and G. Rozenberg.

Theory of traces. Theoretical Computer Science, 60(1):1–82, March 1988. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tiuryn:1988:SRB

[459] J. Tiuryn and P. Urzyczyn. Some relationships between logics of programs and complexity theory. Theoretical Computer Science, 60(1):83–108, March 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

America:1988:DES

[460] P. America and J. De Bakker. Designing equivalent semantic models for process creation. Theoretical Computer Science, 60(2):109–176, September 1988. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Roscoe:1988:LOP

[461] A. W. Roscoe and C. A. R. Hoare. The laws of occam programming. Theoretical Computer Science, 60(2):177–229, September 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Eilam-Tzoreff:1988:MPS

[462] Tali Eilam-Tzoreff and Uzi Vishkin. Matching patterns in strings subject to multi-linear transformations. Theoretical Computer Science, 60(3):231– 254, December 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Duval:1988:GSC

[463] Jean-Pierre Duval. Génération d’une section des classes de conjugaison et arbre des mots de Lyndon de longueur bornée. (French) [Generation of a section of conjugation classes and tree of Lyndon words of limited length]. Theoretical Computer Science, 60(3):255– 283, December 1988. CODEN TCSCDI.

ISSN 0304-3975 (print), 1879-2294 (electronic).

Carpi:1988:SUA

[464] Arturo Carpi. On synchronizing unambiguous automata. Theoretical Computer Science, 60(3):285–296, December 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beeson:1988:TCS

[465] Michael J. Beeson. Towards a computation system based on set theory. Theoretical Computer Science, 60(3):297– 340, December 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Spehner:1988:RFL

[466] Jean-Claude Spehner. La reconnaissance des facteurs d’un langage fini dans un texte en temps linéaire. (French) [The recognition of the factors of a finite language in a text in linear time]. Theoretical Computer Science, 60(3):341– 381, December 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shallit:1988:GAS

[467] J. Shallit. A generalization of automatic sequences. Theoretical Computer Science, 61(1):1–16, October 1988. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Domosi:1988:CCP

[468] P. Domosi and Z. Esik. Critical classes for the α_{0}-product. Theoretical Computer Science, 61(1):17–24, October 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Meznik:1988:SIR

[469] I. Meznik. On a subclass of infinity -regular languages. Theoretical Computer Science, 61(1):25–32, October 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

He:1988:NOP

[470] Xin He. A nearly optimal parallel algorithm for constructing maximal independent set in planar graphs. Theoretical Computer Science, 61(1):33–47, October 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Seger:1988:OTS

[471] C.-J. Seger and J. A. Brzozowski. An optimistic ternary simulation of gate races. Theoretical Computer Science, 61(1):49– 66, October 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vandenBroek:1988:CTG

[472] P. M. van den Broek. Comparison of two graph-rewrite systems. Theoretical Computer Science, 61(1):67–81, October 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Thatte:1988:IFO

[473] S. Thatte. Implementing first-order rewriting with constructor systems. Theoretical Computer Science, 61(1):83– 92, October 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cori:1988:AA

[474] R. Cori, E. Sopena, M. Latteux, and Y. Roos. 2-asynchronous automata.

Theoretical Computer Science, 61(1):93– 102, October 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Book:1988:SGC

[475] Ronald V. Book and Ding-Zhu Du.

The structure of generalized complexity cores. Theoretical Computer Science, 61 (2–3):103–119, November 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dahlhaus:1988:PCP

[476] Elias Dahlhaus and Marek Karpinski.

Parallel construction of perfect matchings and Hamiltonian cycles on dense graphs. Theoretical Computer Science, 61(2–3):121–136, November 1988. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Moriya:1988:ACA

[477] Tetsuo Moriya and Hideki Yamasaki.

Accepting conditions for automata on ω-languages. Theoretical Computer Science, 61(2–3):137–147, November 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

King:1988:AMF

[478] K. N. King. Alternating multihead finite automata. Theoretical Computer Science, 61(2–3):149–174, November 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DiBattista:1988:APR

[479] Giuseppe Di Battista and Roberto Tamassia. Algorithms for plane representations of acyclic digraphs. Theoretical Computer Science, 61(2–3):175– 198, November 1988. CODEN TCSCDI.

ISSN 0304-3975 (print), 1879-2294 (electronic).

Gischer:1988:ETP

[480] Jay L. Gischer. The equational theory of pomsets. Theoretical Computer Science, 61(2–3):199–224, November 1988. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alexi:1988:EVP

[481] Werner Alexi. Extraction and verification of programs by analysis of formal proofs. Theoretical Computer Science, 61(2–3):225–258, November 1988. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gargov:1988:DLC

[482] George Gargov and Solomon Passy. Determinism and looping in combinatory PDL. Theoretical Computer Science, 61 (2–3):259–277, November 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Staiger:1988:SEU

[483] Ludwig Staiger. Ein Satz über die Entropie von Untermonoiden. (German) [A theorem on the entropy of submonoids]. Theoretical Computer Science, 61(2–3): 279–282, November 1988. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kretschmer:1988:CPR

[484] Thomas Kretschmer. A closure property of regular languages. Theoretical Computer Science, 61(2–3):283–287, November 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Arnold:1988:LDF

[485] Andre Arnold. Logical definability of fixed points. Theoretical Computer Science, 61(2–3):289–297, November 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Belaga:1988:BMR

[486] Edward G. Belaga. Bilinear mincing rank. Theoretical Computer Science, 61 (2–3):299–306, November 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Megiddo:1988:FMD

[487] Nimrod Megiddo and Uzi Vishkin. On finding a minimum dominating set in a tournament. Theoretical Computer Science, 61(2–3):307–316, November 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kennaway:1988:GR

[488] R. Kennaway. On ’on graph rewritings’.

Theoretical Computer Science, 61(2–3): 317–320, November 1988. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abiteboul:1988:RHD

[489] Serge Abiteboul and R. Hull. Restructuring hierarchical database objects. Theoretical Computer Science, 62 (1–2):3–38, December 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Atzeni:1988:SCI

[490] Paolo Atzeni and D. Stott Parker. Set containment inference and syllogisms. Theoretical Computer Science, 62(1–2):

39–65, December 1988. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chan:1988:DAB

[491] Edward P. F. Chan and Hector J. Hernandez. On the desirability of γ-acyclic BCNF database schemes. Theoretical Computer Science, 62(1–2):67–104, December 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lakshmanan:1988:SFM

[492] V. S. Lakshmanan. Split-freedom and MVD-intersection: A new characterization of multivalued dependencies having conflict-free covers. Theoretical Computer Science, 62(1–2):105–122, December 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lynch:1988:ITN

[493] Nancy Lynch and Michael Merritt. Introduction to the theory of nested transactions. Theoretical Computer Science, 62(1–2):123–185, December 1988. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sacca:1988:GCM

[494] Domenico Sacca and Carlo Zaniolo. The generalized counting method for recursive logic queries. Theoretical Computer Science, 62(1–2):187–220, December 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vanGucht:1988:IFM

[495] Dirk van Gucht. Interaction-free multivalued dependency sets. Theoretical Computer Science, 62(1–2):221–233, December 1988. CODEN TCSCDI.

ISSN 0304-3975 (print), 1879-2294 (electronic).

Geffert:1988:RRE

[496] Viliam Geffert. A representation of a recursively enumerable languages by two homomorphisms and a quotient. Theoretical Computer Science, 62(3):235– 249, December 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rao:1988:ACA

[497] Nageswara S. V. Rao, S. S. Iyengar, and R. L. Kashyap. An average-case analysis of MAT and inverted file. Theoretical Computer Science, 62(3):251– 266, December 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Salomaa:1988:PRC

[498] Kai Salomaa. A pumping result for 2-context-free languages. Theoretical Computer Science, 62(3):267–287, December 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zeugmann:1988:PRO

[499] Thomas Zeugmann. On the power of recursive optimizers. Theoretical Computer Science, 62(3):289–310, December 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Buss:1988:RPG

[500] Samuel R. Buss and Gyorgy Turan. Resolution proofs on generalized pigeonhole principles. Theoretical Computer Science, 62(3):311–317, December 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Meinel:1988:PNP

[501] Christoph Meinel. The power of nondeterminism in polynomial-size boundedwidth branching programs. Theoretical Computer Science, 62(3):319–325, December 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schmidt:1989:APT

[502] Ursula Schmidt. Avoidable patterns on two letters. Theoretical Computer Science, 63(1):1–17, January 1989. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Livesey:1989:SFB

[503] M. Livesey. Stable families of behavioural equivalences. Theoretical Computer Science, 63(1):19–41, January 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ambos-Spies:1989:RCH

[504] Klaus Ambos-Spies. On the relative complexity of hard problems for complexity classes without complete problems. Theoretical Computer Science, 63 (1):43–61, January 1989. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ginsburg:1989:COH

[505] Seymour Ginsburg and Chang jie Tang. Cohesion of object histories. Theoretical Computer Science, 63(1):63–90, January 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Johnson:1989:UFD

[506] J. Howard Johnson. A unified framework for disambiguating finite transductions.

Theoretical Computer Science, 63(1):91– 111, January 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hagiya:1989:GPP

[507] Masami Hagiya. Generalization from partial parametrization in higher-order type theory. Theoretical Computer Science, 63(2):113–139, February 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Birget:1989:CIT

[508] Jean-Camille Birget. Concatenation of inputs in a two-way automaton. Theoretical Computer Science, 63(2):141– 156, February 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deFelice:1989:CFF

[509] Clelia de Felice. Construction of a family of finite maximal codes. Theoretical Computer Science, 63(2):157– 184, February 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pelc:1989:SKE

[510] Andrzej Pelc. Searching with known error probability. Theoretical Computer Science, 63(2):185–202, February 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hromkovic:1989:TLR

[511] Juraj Hromkovic. Tradeoffs for language recognition on alternating machines. Theoretical Computer Science, 63(2):203–221, February 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Saito:1989:SBE

[512] Takashi Saito and Hidenosuke Nishio. Structural and behavioral equivalence relations in automata networks. Theoretical Computer Science, 63(2):223– 237, February 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Du:1989:ISC

[513] Ding-Zhu Du and Ronald V. Book. On inefficient special cases of NP-complete problems. Theoretical Computer Science, 63(3):239–252, March 1989. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gardarin:1989:TLP

[514] Georges Gardarin, Irene Guessarian, and Christophe de Maindreville. Translation of logic programs into functional fixpoint equations. Theoretical Computer Science, 63(3):253–274, March 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Benson:1989:FPF

[515] David B. Benson and Jerzy Tiuryn. Fixed points in free process algebras, part I. Theoretical Computer Science, 63 (3):275–294, March 1989. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lingas:1989:SIB

[516] Andrzej Lingas. Subgraph isomorphism for biconnected outerplanar graphs in cubic time. Theoretical Computer Science, 63(3):295–302, March 1989. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bloom:1989:ELC

[517] Stephen L. Bloom and Zoltan Esik.

Equational logic of circular data type specification. Theoretical Computer Science, 63(3):303–331, March 1989. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deLuca:1989:SCP

[518] Aldo de Luca and Stefano Varricchio.

Some combinatorial properties of the Thue–Morse sequence and a problem in semigroups. Theoretical Computer Science, 63(3):333–348, March 1989. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stoss:1989:RRF

[519] Hans-Jörg Stoss. On the representation of rational functions of bounded complexity. Theoretical Computer Science, 64(1):1–13, April 1989. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stoss:1989:LBC

[520] Hans-Jörg Stoss. Lower bounds for the complexity of polynomials. Theoretical Computer Science, 64(1):15–23, April 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jojczyk:1989:IBP

[521] K. Jojczyk, J. Konieczny, and T. Kuzak.

On interleaving behaviour of PT-nets. Theoretical Computer Science, 64(1):25– 38, April 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tomita:1989:DBA

[522] Etsuji Tomita and Kazushi Seino. A direct branching algorithm for checking the equivalence of two deterministic pushdown transducers, one of which is real-time strict. Theoretical Computer Science, 64(1):39–53, April 1989. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sato:1989:CBC

[523] Hiroyuki Sato. E-CCC: between CCC and TOPOS. Theoretical Computer Science, 64(1):55–66, April 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hung:1989:SCP

[524] Dang Van Hung and Elod Knuth. Semicommutations and Petri nets. Theoretical Computer Science, 64(1):67– 81, April 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shamir:1989:CAN

[525] Eli Shamir and Assaf Schuster. Communication aspects of networks based on geometric incidence relations. Theoretical Computer Science, 64(1):83–96, April 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hindley:1989:BCL

[526] J. Roger Hindley. BCK-combinators and linear $lambda@-terms have types. Theoretical Computer Science, 64(1):97– 105, April 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Galil:1989:SDP

[527] Zvi Galil and Raffaele Giancarlo. Speeding up dynamic programming with applications to molecular biology. Theoretical Computer Science, 64(1):107– 118, April 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Protasi:1989:NAO

[528] Marco Protasi and Maurizio Talamo. On the number of arithmetical operations for finding Fibonacci numbers. Theoretical Computer Science, 64(1):119– 124, April 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Korach:1989:OLB

[529] E. Korach, S. Moran, and S. Zaks. Optimal lower bounds for some distributed algorithms for a complete network of processors. Theoretical Computer Science, 64(1):125–132, April 1989. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kruskal:1989:TPM

[530] Clyde P. Kruskal, Larry Rudolph, and Marc Snir. Techniques for parallel manipulation of sparse matrices. Theoretical Computer Science, 64(2):135– 157, May 7, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Robert:1989:OSA

[531] Yves Robert and Denis Trystram. Optimal scheduling algorithms for parallel Gaussian elimination. Theoretical Computer Science, 64(2):159–173, May 7, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lyon:1989:DFP

[532] Gordon Lyon. Design factors for parallel processing benchmarks. Theoretical Computer Science, 64(2):175–189, May 7, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bermond:1989:ICE

[533] J. C. Bermond and J. M. Fourneau. Independent connections: An easy characterization of baseline-equivalent multistage interconnection networks. Theoretical Computer Science, 64(2):191– 201, May 7, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Akyildiz:1989:ESO

[534] I. F. Akyildiz and H. von Brand. Exact solutions for open, closed and mixed queueing networks with rejection blocking. Theoretical Computer Science, 64 (2):203–219, May 7, 1989. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stark:1989:CTS

[535] Eugene W. Stark. Concurrent transition systems. Theoretical Computer Science, 64(3):221–269, May 29, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Therien:1989:PAM

[536] Denis Therien. Programs over aperiodic monoids. Theoretical Computer Science, 64(3):271–280, May 29, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mazurkiewicz:1989:CSI

[537] Antoni Mazurkiewicz, Edward Ochmanski, and Wojciech Penczek. Concurrent systems and inevitability. Theoretical Computer Science, 64(3):281–304, May 29, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Howell:1989:PCF

[538] Rodney R. Howell and Louis E. Rosier.

Problems concerning fairness and temporal logic for conflict-free Petri nets. Theoretical Computer Science, 64(3): 305–329, May 29, 1989. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alon:1989:NTB

[539] Noga Alon and Uri Zwick. On Neciporuk’s theorem for branching programs. Theoretical Computer Science, 64(3):331–342, May 29, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

KleineBuning:1989:PVA

[540] Hans Kleine Büning, Theodor Lettmann, and Ernst W. Mayr. Projections of vector addition system reachability sets are semilinear. Theoretical Computer Science, 64(3):343–350, May 29, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Enjalbert:1989:MRC

[541] Patrice Enjalbert and Luis Farinas del Cerro. Modal resolution in clausal form. Theoretical Computer Science, 65(1):1– 33, June 15, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abadi:1989:PTP

[542] Martin Abadi. The power of temporal proofs. Theoretical Computer Science, 65(1):35–83, June 15, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Thatte:1989:FAL

[543] Satish R. Thatte. Full abstraction and limiting completeness in equational languages. Theoretical Computer Science, 65(1):85–119, June 15, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:1989:CAC

[544] Anonymous. Conference on Arithmetics and Coding Systems. Theoretical Computer Science, 65(2):??, June 28, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Allouche:1989:SRF

[545] Jean-Paul Allouche. On a sequence of rational functions. Theoretical Computer Science, 65(2):123–130, June 28, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blanchard:1989:BES

[546] F. Blanchard. $beta@-expansions and symbolic dynamics. Theoretical Computer Science, 65(2):131–141, June 28, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bleuzen-Guernalec:1989:PCR

[547] Noelle Bleuzen-Guernalec. On a possible classification of real-time constructed sequences. Theoretical Computer Science, 65(2):143–148, June 28, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dekking:1989:POL

[548] F. M. Dekking. On the probability of occurrence of labelled subtrees of a randomly labelled tree. Theoretical Computer Science, 65(2):149–152, June 28, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dumont:1989:SNR

[549] J.-M. Dumont and Alain Thomas.

Systémes de numération et fonction fractales relatifs aux substitutions. (French) [Systems of numeration and fractal functions relative to substitutions]. Theoretical Computer Science, 65(2):153– 169, June 28, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hansel:1989:RPM

[550] G. Hansel and D. Perrin. Rational probability measures. Theoretical Computer Science, 65(2):171–188, June 28, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hellekalek:1989:WSS

[551] P. Hellekalek and G. Larcher. On Weyl sums and skew products over irrational rotations. Theoretical Computer Science, 65(2):189–196, June 28, 1989. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kraaikamp:1989:SEP

[552] Cor Kraaikamp. Statistic and ergodic properties of Minkowski’s diagonal continued fraction. Theoretical Computer Science, 65(2):197–212, June 28, 1989.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

MendesFrance:1989:GEI

[553] M. Mendes France and A. J. van der Poorten. From geometry to Euler identities. Theoretical Computer Science, 65 (2):213–220, June 28, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mignosi:1989:IWL

[554] Filippo Mignosi. Infinite words with linear subword complexity. Theoretical Computer Science, 65(2):221–242, June 28, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mori:1989:FDP

[555] Makoto Mori. On the Fredholm determinant of a piecewise linear transformation. Theoretical Computer Science, 65 (2):243–248, June 28, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mosse:1989:QAS

[556] Brigitte Mosse. Q-adic spectral analysis of some arithmetic sequences. Theoretical Computer Science, 65(2):249– 263, June 28, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Restivo:1989:FGS

[557] Antonio Restivo. Finitely generated sofic systems. Theoretical Computer Science, 65(2):265–270, June 28, 1989. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yokouchi:1989:CRT

[558] Hirofumi Yokouchi. Church–Rosser theorem for a rewriting system on categorical combinators. Theoretical Computer Science, 65(3):271–290, July 17, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hardin:1989:CRP

[559] Therese Hardin. Confluence results for the pure strong categorical logic CCL $lambda@-calculi as subsystems of CCL. Theoretical Computer Science, 65(3): 291–342, July 17, 1989. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shepherdson:1989:SCS

[560] J. C. Shepherdson. A sound and complete semantics for a version of negation as failure. Theoretical Computer Science, 65(3):343–371, July 17, 1989. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lehmkuhl:1989:OAA

[561] T. Lehmkuhl and T. Lickteig. On the order of approximation in approximative triadic decompositions of tensors. Theoretical Computer Science, 66(1):1– 14, August 2, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dunne:1989:MSN

[562] P. E. Dunne. On monotone simulations of nonmonotone networks. Theoretical Computer Science, 66(1):15–25, August 2, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Piperno:1989:APC

[563] A. Piperno. Abstraction problems in combinatory logic: a compositive approach. Theoretical Computer Science, 66(1):27–43, August 2, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kari:1989:OCP

[564] J. Kari. Observations concerning a public-key cryptosystem based on iterated morphisms. Theoretical Computer Science, 66(1):45–53, August 2, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Xin:1989:EAD

[565] Zhang Luo Xin. An efficient algorithm to decide whether a monoid presented by a regular Church–Rosser Thue system is a group. Theoretical Computer Science, 66(1):55–63, August 2, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

opdenAkker:1989:LGL

[566] R. op den Akker. On LC(0) grammars and languages. Theoretical Computer Science, 66(1):65–85, August 2, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Urquhart:1989:CGS

[567] A. Urquhart. The complexity of Gentzen systems for propositional logic. Theoretical Computer Science, 66(1):87–97, August 2, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Statman:1989:SSC

[568] R. Statman. On sets of solutions to combinator equations. Theoretical Computer Science, 66(1):99–104, August 2, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pelc:1989:WAC

[569] A. Pelc. Weakly adaptive comparison searching. Theoretical Computer Science, 66(1):105–111, August 2, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mundici:1989:FCM

[570] D. Mundici. Functions computed by monotone Boolean formulas with no repeated variables. Theoretical Computer Science, 66(1):113–114, August 2, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diekert:1989:KBC

[571] Volker Diekert. On the Knuth–Bendix completion for concurrent processes. Theoretical Computer Science, 66(2): 117–136, August 20, 1989. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dietzfelbinger:1989:LBS

[572] Martin Dietzfelbinger. Lower bounds for sorting of sums. Theoretical Computer Science, 66(2):137–155, August 20, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Edelsbrunner:1989:TNC

[573] Herbert Edelsbrunner, Guenter Rote, and Ermo Welzl. Testing the necklace condition for shortest tours and optimal factors in the plane. Theoretical Computer Science, 66(2):157–180, August 20, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Levcopoulos:1989:HOB

[574] Christos Levcopoulos, Andrzej Lingas, and Jorg-R. R. Sack. Heuristics for optimum binary search trees and minimum weight triangulation problems. Theoretical Computer Science, 66(2):181–203, August 20, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

NaitAbdallah:1989:LAA

[575] M. A. Nait Abdallah. A logico-algebraic approach to the model theory of knowledge. Theoretical Computer Science, 66 (2):205–232, August 20, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Weil:1989:IMD

[576] P. Weil. Inverse monoids of dot-depth two. Theoretical Computer Science, 66 (3):233–245, August 26, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yamasaki:1989:LTR

[577] H. Yamasaki. Language-theoretical representations of gw-languages. Theoretical Computer Science, 66(3):247–254, August 26, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Angluin:1989:TS

[578] D. Angluin, W. I. Gasarch, and C. H. Smith. Training sequences. Theoretical Computer Science, 66(3):255–272, August 26, 1989. CODEN TCSCDI.

ISSN 0304-3975 (print), 1879-2294 (electronic).

Ito:1989:DTD

[579] A. Ito, K. Inoue, and I. Takanami. Deterministic two-dimensional on-line tessellation acceptors are equivalent to twoway two-dimensional alternating finite automata through 180 degrees-rotation. Theoretical Computer Science, 66(3): 273–287, August 26, 1989. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jacopini:1989:GIF

[580] G. Jacopini and P. Mentrasti. Generation of invertible functions. Theoretical Computer Science, 66(3):289–297, August 26, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Makowsky:1989:WSO

[581] J. A. Makowsky and I. Sain. Weak second order characterizations of various program verification systems. Theoretical Computer Science, 66(3):299–321, August 26, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mezghiche:1989:PCB

[582] M. Mezghiche. On pseudo-c beta normal form in combinatory logic. Theoretical Computer Science, 66(3):323–331, August 26, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tiomkin:1989:PTV

[583] M. Tiomkin. Probabilistic termination versus fair termination. Theoretical Computer Science, 66(3):333–340, August 26, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Honkala:1989:NCR

[584] J. Honkala. A necessary condition for the rationality of the zeta function of a regular language. Theoretical Computer Science, 66(3):341–347, August 26, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Swart:1989:IS

[585] Charles Swart and Dana Richards. On the inference of strategies. Theoretical Computer Science, 67(1):5–18, September 5, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Otto:1989:DCF

[586] Friedrich Otto. On deciding confluence of finite string-rewriting systems modulo partial commutativity. Theoretical Computer Science, 67(1):19–35, September 5, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Martin:1989:GAM

[587] Ursula Martin. A geometrical approach to multiset orderings. Theoretical Computer Science, 67(1):37–54, September 5, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Clausen:1989:FGF

[588] Michael Clausen. Fast generalized Fourier transforms. Theoretical Computer Science, 67(1):55–63, September 5, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beauquier:1989:MAF

[589] Daniele Beauquier. Minimal automation for a factorial, transitive, and rational language. Theoretical Computer Science, 67(1):65–73, September 5, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Moriya:1989:GCA

[590] Etsuro Moriya. A grammatical characterization of alternating pushdown automata. Theoretical Computer Science, 67(1):75–85, September 5, 1989. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Marongiu:1989:CBT

[591] G. Marongiu and S. Tulipani. On a conjecture of Bergstra and Tucker. Theoretical Computer Science, 67(1):87–97, September 5, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Inoue:1989:LSH

[592] Katsushi Inoue, Itsuo Takanami, and Juraj Hromkovic. Leaf-size hierarchy of two-dimensional alternating Turing machines. Theoretical Computer Science, 67(1):99–110, September 5, 1989. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pardubska:1989:NMM

[593] Dana Pardubska and Ivana Stefanekova.

Nondeterministic multicounter machines and complementation. Theoretical Computer Science, 67(1):111–113, September 5, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cosnard:1989:CSX

[594] M. Cosnard, J. Duprat, and A. Ferreira. Complexity of selection in X + Y. Theoretical Computer Science, 67(1): 115–120, September 5, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Petuaud:1989:ETS

[595] Catherine Petuaud. Entropie topologique des systèmes specifiés. (French) [Topological entropy of specified systems]. Theoretical Computer Science, 67 (1):121–128, September 5, 1989. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

PinoPerez:1989:DRE

[596] Ramon Pino Perez. Decidability of the restriction equational theory in the partial lambda calculus. Theoretical Computer Science, 67(1):129–139, September 5, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Madlener:1989:ADP

[597] K. Madlener and F. Otto. About the descriptive power of certain classes of finite string-rewriting systems. Theoretical Computer Science, 67(2–3):143– 172, October 3, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bachmair:1989:CRM

[598] L. Bachmair and N. Dershowitz. Completion for rewriting modulo a congruence. Theoretical Computer Science, 67 (2–3):173–201, October 3, 1989. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gallier:1989:CST

[599] J. H. Gallier and W. Snyder. Complete sets of transformations for general E-unification. Theoretical Computer Science, 67(2–3):203–260, October 3, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Choppy:1989:CAT

[600] C. Choppy, S. Kaplan, and M. Soria.

Complexity analysis of term-rewriting systems. Theoretical Computer Science, 67(2–3):261–282, October 3, 1989. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baeten:1989:TRS

[601] J. C. M. Baeten, J. A. Bergstra, J. W.

Klop, and W. P. Weijland. Termrewriting systems with rule priorities. Theoretical Computer Science, 67(2–3): 283–301, October 3, 1989. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kirchner:1989:SIS

[602] H. Kirchner. Schematization of infinite sets of rewrite rules generated by divergent completion processes. Theoretical Computer Science, 67(2–3):303–332, October 3, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kirschenhofer:1989:BPP

[603] P. Kirschenhofer, H. Prodinger, and W. Szpankowski. On the balance property of Patricia trees: external path length viewpoint. Theoretical Computer Science, 68(1):1–17, October 16, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chang:1989:ESS

[604] J. H. Chang, O. H. Ibarra, and M. A. Palis. Efficient simulations of simple models of parallel computation by time-bounded ATMs and space-bounded TMs. Theoretical Computer Science, 68 (1):19–36, October 16, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Droste:1989:ESD

[605] M. Droste. Event structures and domains. Theoretical Computer Science, 68(1):37–47, October 16, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hofer:1989:LIR

[606] G. Hofer. Left ideals and reachability in machines. Theoretical Computer Science, 68(1):49–56, October 16, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gambosi:1989:WCA

[607] G. Gambosi, G. F. Italiano, and M. Talamo. Worst-case analysis of the setunion problem with extended backtracking. Theoretical Computer Science, 68 (1):57–70, October 16, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Huckenbeck:1989:EGT

[608] U. Huckenbeck. Euclidian geometry in terms of automata theory. Theoretical Computer Science, 68(1):71–87, October 16, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Karloff:1989:NAB

[609] H. J. Karloff. An NC algorithm for Brooks’ theorem. Theoretical Computer Science, 68(1):89–103, October 16, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Josephs:1989:SLF

[610] M. B. Josephs. The semantics of lazy functional languages. Theoretical Computer Science, 68(1):105–111, October 16, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chlebus:1989:HPH

[611] B. S. Chlebus. A hierarchy of propositional Horn formulas. Theoretical Computer Science, 68(1):113–119, October 16, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Arvind:1989:SBR

[612] V. Arvind and S. Biswas. On some bandwidth restricted versions of the satisfiability problem of propositional CNF formulas. Theoretical Computer Science, 68 (2):123–134, October 30, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blair:1989:PLP

[613] H. A. Blair and V. S. Subrahmanian.

Paraconsistent logic programming. Theoretical Computer Science, 68(2):135– 154, October 30, 1989. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lingas:1989:PCS

[614] A. Lingas and A. Proskurowski. On parallel complexity of the subgraph homeomorphism and the subgraph isomorphism problem for classes of planar graphs. Theoretical Computer Science, 68(2):155–173, October 30, 1989. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Parrow:1989:SCE

[615] J. Parrow. Submodule construction as equation solving in CCS. Theoretical Computer Science, 68(2):175–202, October 30, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ramanujam:1989:SDD

[616] R. Ramanujam. Semantics of distributed definite clause programs. Theoretical Computer Science, 68(2):203– 220, October 30, 1989. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Coquand:1989:CE

[617] T. Coquand. Categories of embeddings. Theoretical Computer Science, 68 (3):221–237, November 12, 1989. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gutbrod:1989:TSG

[618] R. Gutbrod. A transformation system for generating description languages of chain code pictures. Theoretical Computer Science, 68(3):239–252, November 12, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blanchard:1989:CSS

[619] F. Blanchard. Certain sofic systems engendered codes. Theoretical Computer Science, 68(3):253–265, November 12, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Immerman:1989:RRC

[620] N. Immerman and S. R. Mahaney. Relativizing relativized computations. Theoretical Computer Science, 68(3):267– 276, November 12, 1989. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hortala-Gonzalez:1989:HLN

[621] M. T. Hortala-Gonzalez and M. Rodriguez-Artalejo. Hoare’s logic for nondeterministic regular programs: a nonstandard approach. Theoretical Computer Science, 68(3):277–302, November 12, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bracho:1989:CGF

[622] F. Bracho. Continuously generated fixed points. Theoretical Computer Science, 68(3):303–317, November 12, 1989. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Narendran:1989:SPT

[623] P. Narendran and F. Otto. Some polynomial-time algorithms for finite monadic Church–Rosser Thue systems. Theoretical Computer Science, 68(3): 319–332, November 12, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Solitro:1989:TCB

[624] U. Solitro. A typed calculus based on a fragment of linear logic. Theoretical Computer Science, 68(3):333–342, November 12, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gafni:1989:SEC

[625] E. Gafni, J. Naor, and P. Ragde. On separating the EREW and CREW PRAM models. Theoretical Computer Science, 68(3):343–346, November 12, 1989. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rupert:1989:CRF

[626] C. P. Rupert. Comment on a remark of forys. Theoretical Computer Science, 68 (3):347–348, November 12, 1989. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vieille:1989:RQP

[627] L. Vieille. Recursive query processing: the power of logic. Theoretical Computer Science, 69(1):1–53, December 5, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kerisit:1989:RAL

[628] J.-M. Kerisit. A relational approach to logic programming: the extended Alexander method. Theoretical Computer Science, 69(1):55–68, December 5, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kaplan:1989:ASC

[629] S. Kaplan. Algebraic specification of concurrent systems. Theoretical Computer Science, 69(1):69–115, December 5, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nielson:1989:TLS

[630] F. Nielson. Two-level semantics and abstract interpretation. Theoretical Computer Science, 69(2):117–242, December 11, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Felleisen:1989:STS

[631] M. Felleisen and D. P. Friedman. A syntactic theory of sequential state. Theoretical Computer Science, 69(3):243– 287, December 18, 1989. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Falaschi:1989:DMO

[632] M. Falaschi, G. Levi, C. Palamidessi, and M. Martelli. Declarative modeling of the operational behavior of logic languages. Theoretical Computer Science, 69(3):289–318, December 18, 1989. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baker:1989:GPA

[633] K. A. Baker, G. F. McNulty, and W. Taylor. Growth problems for avoidable words. Theoretical Computer Science, 69(3):319–345, December 18, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chrobak:2003:EFA

[634] Marek Chrobak. Errata to: “Finite Automata and Unary Languages”: [Theoret. Comput. Sci. 47 (1986) 149–158]. Theoretical Computer Science, 302(1–3): 497–498, June 13, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See [240].

Gruska:1988:ISM

[635] Josef Gruska, editor. International symposium on mathematical foundations of computer science, MFCS ’86, volume 57(1) of Theoretical Computer Science. North-Holland, Amsterdam, The Netherlands, 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kott:1988:TIC

[636] Laurent Kott, editor. Thirteenth international colloquium on automata, languages and programming, volume 58(1–3) of Theoretical Computer Science. North-Holland, Amsterdam, The Netherlands, June 1988. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Levi:1988:IJC

[637] Georgio Levi and Ugo Montanari, editors. International joint conference on theory and practice of software development — TAPSOFT ’87, volume 59(1–2) of Theoretical Computer Science. North-Holland, Amsterdam, The Netherlands, July 1988. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ausiello:1988:SPP

[638] Giorgio Ausiello, editor. Selected papers presented at the first international conference on database theory, volume 62(1–2) of Theoretical Computer Science. North-Holland, Amsterdam, The Netherlands, December 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liardet:1989:CAC

[639] P. Liardet and G. Rauzy, editors. Conference on Arithmetics and Coding Systems, volume 65(2) of Theoretical Computer Science. North-Holland, Amsterdam, The Netherlands, June 1989. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ottmann:1989:ICA

[640] Thomas Ottmann, editor. 14th International Colloquium on Automata, Languages and Programming, volume 66(2) of Theoretical Computer Science. North-Holland, Amsterdam, The Netherlands, August 1989. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).